### v2

# A mathematical theory of topological invariants of quantum spin systems

Abstract
We show that Hall conductance and its non-abelian and higher-dimensional analogs are obstructions to promoting a symmetry of a state to a gauge symmetry. To do this, we define a local Lie algebra over a Grothendieck site as a pre-cosheaf of Lie algebras with additional properties and propose that a gauge symmetry should be described by such an object. We show that infinitesimal symmetries of a gapped state of a quantum spin system form a local Lie algebra over a site of semilinear sets and use it to construct topological invariants of the state. Our construction applies to lattice systems on arbitrary asymptotically conical subsets of a Euclidean space including those which cannot be studied using field theory.
1 Introduction
The study of gapped phases of quantum matter at zero temperature is an important area of theoretical physics. Much conceptual progress has been made by assuming that gapped phases can be described by topological quantum field theory (TQFT). For example, the celebrated Quantum Hall Effect is captured by Chern-Simons field theory. However, the precise relation between gapped phases of matter and TQFTs is not understood. Recently, new mathematically rigorous approaches to classifying gapped phases of matter have been developed (see [1, 2, 3, 4] for the case of one-dimensional systems, [5, 6, 7, 8, 9, 10] for the case of two-dimensional systems, and [11] for systems in an arbitrary number of dimensions). They enable one to assign indices to gapped states of infinite-volume quantum systems invariant under symmetries. The main property of these indices, also referred to as topological invariants, is that they do not vary along suitably-defined continuous paths in the space of states. In some cases, the indices can be related to physical quantities, such as the zero-temperature Hall conductance, thereby explaining the robustness of the latter.
The methods of [7, 11, 8, 9] apply to arbitrary gapped states of infinite-volume quantum spin systems with rapidly decaying interactions and employ $C^{*}$-algebraic techniques, some well-established and some relatively new. The construction of topological invariants in [11, 8] also uses some algebraic and geometric ingredients . The algebraic ingredient is a pointed (or curved) Differential Graded Lie Algebra (DGLA) and an associated Maurer-Cartan equation. The geometric ingredient is a collection of conical subsets of the Euclidean space triangulating the sphere at infinity. The appearance of these ingredients in the context of quantum statistical mechanics has not been motivated, and consequently the mathematical meaning of the invariants remains obscure.
The primary goal of this paper is provide a proper mathematical framework for the constructions of [11, 8] and to interpret topological invariants of gapped states as lattice analogs of ’t Hooft anomalies in Quantum Field Theory. The secondary goal is to generalize the construction in various directions. In particular, we show how to define topological invariants of lattice spin systems confined to well-behaved subsets of the lattice. This generalization makes explicit that the invariants take values in a vector space which is determined by the asymptotic geometry of the subset.
While our work concerns quantum lattice systems, we take inspiration from Quantum Field Theory (QFT). These two subjects are connected via the bulk-boundary correspondence. One aspect of this conjectural correspondence is that topological invariants of gapped states with symmetries are related to ’t Hooft anomalies of symmetries of the boundary field theory. [^1] This assumes, of course, that a field-theoretic description of boundary degrees of freedom exists, which is far from obvious. It is usually said that ’t Hooft anomalies are obstructions to gauging a global symmetry of a QFT [12]. A possible mathematical interpretation of this statement is that an ’t Hooft anomaly is an obstruction to defining a local action of the group of gauge transformations on the algebra of local observables of a QFT. Assuming this interpretation, the presence of an ’t Hooft anomaly is a purely kinematic statement which involves neither the Hamiltonian nor the vacuum state of the field theory. It is not clear if conventional markers of ’t Hooft anomalies, such as anomalous Ward identities for vacuum correlators of currents, are implied by a kinematic statement. Proving or disproving this is currently out of reach because of gaps in the mathematical foundations of QFT. The mathematical theory of quantum lattice systems, on the other hand, is sufficiently mature and enables us to address the problem of ’t Hooft anomalies from the bulk side of the bulk-boundary correspondence. In this paper we show that topological invariants of gapped states of lattice systems, such as the zero-temperature Hall conductance, can be interpreted as obstructions to promoting a symmetry of a gapped state to a gauge symmetry. Dynamics enters this statement only though the state.
The main novelty of the paper is a new formulation of locality on a lattice. Building on ideas introduced in [7, 11, 8], we define for any (possibly unbounded) region of the lattice a space of derivations that are approximately localized on that region. For sufficiently regular regions, we show that these spaces behave as expected under natural operations like the commutator. To combine them into a single geometric object we show that they form a cosheaf on a certain site, i. e. a category with a Grothendieck topology. The utility of Grothendieck topologies in describing spaces of functions with a prescribed asymptotic behavior is well-known to analysts, see e.g. [13, 14], and here we apply the same idea in a non-commutative context. We call the resulting global geometric object a local Lie algebra. An on-site action of a Lie group on a lattice system yields a representation of a certain local Lie algebra by derivations, and we show that the invariants defined in [11] can be phrased as purely algebraic invariants of this representation. Namely, an obstruction exists to defining a representation which acts by state-preserving derivations, and this obstruction takes value in the homology of a certain DGLA associated naturally to the local Lie algebra of state-preserving derivations. The invariants defined in [11] then arise from a natural pairing of this homology with the Čech cohomology of the sphere at infinity. Aside from clarifying their mathematical meaning, this also shows that the invariants defined in [11] do not depend on certain choices present in their construction.
The content of the paper is as follows. In Section 2 we axiomatize the notion of an infinitesimal local symmetry by defining local Lie algebras abstractly in terms of cosheaves on a site. We also introduce the Čech functor which assigns a DGLAs to a local Lie algebra, and will serve as the main link between lattice geometry and algebraic topology. In Section 3 we turn to infinitesimal symmetries on the lattice, introducing the space of derivations approximately localized on a region, and prove the main properties of these spaces, some of which hold only for sufficiently regular regions. In section 4 we identify a suitable category of such regular regions in ${\mathbb{R}}^{n}$, the category ${\mathcal{C}\mathcal{S}}_{n}$ of fuzzy semilinear sets. We identify a Grothendieck topology on ${\mathcal{C}\mathcal{S}}_{n}$ and show that infinitesimal symmetries of any gapped state $\psi$ of a quantum lattice system on ${\mathbb{R}}^{n}$ can be described by a local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. We also study gapped states invariant under an action of a compact Lie group $G$ and define a $G$-equivariant analog of ${\mathfrak{D}}^{\psi}_{al}$. In Section 6 we construct invariants of $G$-invariant gapped states. The construction is along the lines of [11, 8] and uses an inhomogeneous Maurer-Cartan equation. We show that these invariants are obstructions for promoting the symmetry $G$ to a local symmetry of the state $\psi$. We also explain how to generalize the construction of invariants to lattice systems defined on sufficiently nice (asymptotically conical) subsets of ${\mathbb{R}}^{n}$ and show that their invariants take values in a space which depends on the asymptotic geometry of the subset. This goes beyond what one can access using the TQFT heuristics. In Appendix A we isolate some proofs necessary for Section 3, and Appendix B develops some properties of the inhomogeneous Maurer-Cartan equation.
We are grateful to Owen Gwilliam, Ezra Getzler, and Bas Janssens for discussions. The work of A. A. and A. K. was supported by the Simons Investigator Award. B. Y. would like to thank Yu-An Chen and Peking University for hospitality during the final stages of this work. A. K. would like to thank Yau Mathematical Sciences Center, Tsinghua University, for the same.
2 Local Lie algebras
2.1 Locality and (pre-)cosheaves
Let $M$ be a manifold and $Open(M)$ be the category whose objects are open subsets of $M$, and the set of morphisms from an open $U$ to an open $V$ is the singleton or the empty set depending on whether $U\subseteq V$ or $U\nsubseteq V$. Composition of morphisms is uniquely defined. A pre-cosheaf ${\mathfrak{F}}$ on $M$ with values in a category ${\mathcal{C}}$ is a functor ${\mathfrak{F}}:Open(M)\rightarrow{\mathcal{C}}$. Thus for every inclusion of opens $U\subseteq V$ one is given a co-restriction morphism $e_{VU}:{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(V)$ such that for any three opens $U\subseteq V\subseteq W$ one has $e_{WV}\circ e_{VU}=e_{WU}$. Pre-cosheaves (as well as pre-sheaves, which are functors from the opposite category of $Open(M)$ to ${\mathcal{C}}$) can be used to describe local data on $M$. This form of locality is rather weak, since it does not require ${\mathfrak{F}}(U\cup V)$ to be expressible through ${\mathfrak{F}}(U)$ and ${\mathfrak{F}}(V)$.
As an example, consider the Lie algebra of gauge transformations, i.e. the Lie algebra ${\mathfrak{G}}(M):=C^{\infty}(M,{\mathfrak{g}})$ of smooth functions on $M$ with values in a finite-dimensional Lie algebra ${\mathfrak{g}}$. It is a global object attached to $M$. To “localize” it, for any open $U\subseteq M$ we define the Lie algebra ${\mathfrak{G}}(U)$ to be the space of smooth ${\mathfrak{g}}$-valued functions on $M$ whose closed support is contained in $U$. In particular, ${\mathfrak{G}}(\varnothing)=0$. For any inclusion of opens $U\subseteq V$ we have a homomorphism of Lie algebras $\iota_{VU}:{\mathfrak{G}}(U)\rightarrow{\mathfrak{G}}(V)$ such that the Lie algebras ${\mathfrak{G}}(U)$ assemble into a pre-cosheaf ${\mathfrak{G}}$ of Lie algebras on $M$. This is a coflasque pre-cosheaf, ie. all its structure maps $\iota_{VU}$ are injective. [^2] The terminology comes from sheaf theory, where a pre-sheaf ${\mathcal{F}}:Open(M)^{opp}\rightarrow{\mathcal{C}}$ is called flasque if for any $U\subseteq V$ the restriction morphism ${\mathcal{F}}(V)\rightarrow{\mathcal{F}}(U)$ is a surjection.
Continuing with the example, for any two opens $U,V$ the following sequence is exact:
${\mathfrak{G}}(U\cap V)\rightarrow{\mathfrak{G}}(U)\oplus{\mathfrak{G}}(V)\rightarrow{\mathfrak{G}}(U\cup V)\rightarrow 0.$ (1)
Here the first arrow is $\iota_{U,U\cap V}\oplus(-\iota_{V,U\cap V})$ and the second arrow is $\iota_{U\cup V,U}\oplus\iota_{U\cup V,V}$. Exactness follows from the existence of a partition of unity for the cover ${\mathfrak{U}}=\{U,V\}$ of $U\cup V$. In words, the exactness of the sequence (1) means that any element of ${\mathfrak{G}}(U\cup V)$ can be decomposed as a sum of elements of sub-algebras attached to $U$ and $V$ modulo ambiguities which take values in the sub-algebra attached to $U\cap V$.
More generally, the existence of a partition of unity implies that for any collection of opens $U_{i}$, $i\in I$, the following sequence is exact:
$\oplus_{i<j}{\mathfrak{G}}(U_{i}\cap U_{j})\rightarrow\oplus_{i}{\mathfrak{G}}(U_{i})\rightarrow{\mathfrak{G}}(\cup_{i}U_{i})\rightarrow 0.$ (2)
By definition, this means that the pre-cosheaf of vector spaces ${\mathfrak{G}}$ is a cosheaf of vector spaces. The cosheaf property is a compatibility of the pre-cosheaf with the notions of intersection and union of opens and expresses a stronger form of locality.
Note that the maps in the above exact sequences are not Lie algebra homomorphisms. Hence ${\mathfrak{G}}$ is not a cosheaf of Lie algebras. Nevertheless, the following additional property of ${\mathfrak{G}}$ can be regarded as a form of locality of the Lie bracket: for any $U,V,$ $[{\mathfrak{G}}(U),{\mathfrak{G}}(V)]\subseteq{\mathfrak{G}}(U\cap V)$. We will call this ”Property I”, where ”I” stands for ”intersection”. In particular, elements of ${\mathfrak{G}}(M)$ which are supported on non-intersecting opens $U$ and $V$ commute.
Symmetries of gapped states of quantum lattice spin systems are typically local only approximately. To phrase locality of symmetries in lattice systems in a similar language, one needs to replace the set $Open(M)$ of open subsets of $M$ with a more general structure which admits the notions of intersection, union, and cover.
The first thing to note is that the category $Open(M)$ is rather special: its objects form a pre-ordered set (i.e. the set of objects carries a relation $\subseteq$ which is reflexive and transitive), and the category structure is determined by the pre-order. The relation $\subseteq$ is also anti-symmetric: $U\subseteq V$ and $V\subseteq U$ implies $U=V$. In other words, $Open(M)$ is a poset. In general, we will not require the pre-order to be anti-symmetric. From the categorical viewpoint, $U\subseteq V$ and $V\subseteq U$ means that $U$ and $V$ are isomorphic objects of the category $Open(M)$, and as a general rule, it is not advisable to identify isomorphic objects.
For any pre-ordered set $(X,\leq)$ there is a natural notion of intersection and union. The intersection of $U,V\in X$ can be defined as the greatest lower bound (or meet) of both $U$ and $V$, i.e. a $W\in X$ such that $W\leq U$, $W\leq V$, and for any $W^{\prime}\leq U,V$ we have $W^{\prime}\leq W$. The meet of $U$ and $V$ is denoted $U\wedge V$. Similarly, the union of $U$ and $V$ can be defined as the smallest upper bound (or join) of both $U$ and $V$. It is denoted $U\vee V$. For a general pre-ordered set, the meet and join may not exist for all pairs of objects. If they exist, they are unique up to isomorphism. We will assume that $(X,\leq)$ is such that $U\wedge V$ and $U\vee V$ exist for all $U,V\in X$ [^3] If we turn $X$ into a poset by identifying isomorphic objects, then this means that the poset is a lattice in the sense of order theory. The existence of all pairwise meets and joins implies the existence of all finite meets and joins. In the case of the pre-ordered set $Open(M)$ arbitrary (i.e. not necessarily finite) joins make sense.
Finally, to define covers of elements of $X$, let us assume that $U\wedge(V\vee W)\leq(U\wedge V)\vee(U\wedge W)$ for all $U,V,W\in X$ [^4] The opposite relation is automatic, so this condition ensures that $U\wedge(V\vee W)\simeq(U\wedge V)\vee(U\wedge W)$ for all $U,V,W\in X$. This is equivalent to saying that the poset corresponding to $X$ is a distributive lattice. We will say that $X$ is a distributive pre-ordered set. This condition is certainly satisfied for $Open(M)$. We say that a collection ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ of elements of $X$ covers $A\in X$ iff $U_{i}\leq A$ for all $i\in I$ and $A\leq\bigvee_{i\in I}U_{i}$. This definition ensures that if ${\mathfrak{U}}$ covers $A$, then for any $B\leq A$ the collection ${\mathfrak{U}}\wedge B=\{U_{i}\wedge B\}_{i\in I}$ covers $B$.
In the case of the pre-ordered set $Open(M)$, the standard topological definition of a cover allows $I$ to be infinite. In general, if $X$ admits only finite joins, $I$ needs to be finite. Also, we may or may not allow $I$ to be empty. This possibility only arises when $A$ is the smallest element of $X$, i.e. $A\leq U$ for any $U\in X$. In the case of the pre-ordered set $Open(M)$, the smallest element is the empty set $\varnothing$, and the standard choice is to allow the empty cover of the empty set. For a general pre-ordered set, it is up to us whether to allow $I$ to be empty.
From a categorical perspective, this notion of a cover equips any distributive pre-ordered set $(X,\leq)$ admitting pairwise meets and joins with a Grothendieck topology, thus making it into a site [15]. Apart from the option of allowing the labeling set $I$ to be empty, this Grothendieck topology is canonical.
For any $W\in X$ we may consider the subset $X^{W}=\left\{U\in X\mid U\leq W\right\}$ with the pre-order inherited from $(X,\leq)$. It is a distributive pre-ordered set in its own right. When equipped with its canonical Grothendieck topology, it can be regarded as a sub-site of the site associated to $(X,\leq)$.
Given any distributive pre-ordered set $(X,\leq)$, we can define the notion of a pre-cosheaf of vector spaces, a pre-cosheaf of Lie algebras, a cosheaf of vector spaces, and a coflasque pre-cosheaf of Lie algebras with Property I exactly as before, i.e. by mechanically replacing $\cup$ with $\vee$ and $\cap$ with $\wedge$. Motivated by the above example, we introduce the following definition.
Definition 2.1.
A local Lie algebra over $(X,\leq)$ is a coflasque pre-cosheaf of Lie algebras with Property I which is also a cosheaf of vector spaces over the corresponding site. A morphism of local Lie algebras is a morphism of the underlying pre-cosheaves of Lie algebras.
Remark 2.1.
Let ${\mathfrak{F}}$ be a local Lie algebra. Then for any $U\leq V$ the Lie algebra ${\mathfrak{F}}(U)$ is an ideal in ${\mathfrak{F}}(V)$. One can equivalently define a local Lie algebra as a pre-cosheaf of Lie algebras over $(X,\leq)$ which is a cosheaf of vector spaces and such that all co-restriction maps are inclusions of Lie ideals.
Remark 2.2.
In this paper all Lie algebras will be Fréchet-Lie algebras and all morphisms will be continuous. We define a local Fréchet-Lie algebra over $(X,\leq)$ to be a coflasque pre-cosheaf of Fréchet-Lie algebras with Property I which is also a cosheaf of vector spaces.
The following Lemma will be useful in future sections to check the cosheaf property:
Lemma 2.1.
Let ${\mathfrak{F}}$ be a pre-cosheaf of vector spaces on a distributive lattice $X$, with co-restriction maps $\iota_{U,V}:U\to V$. Then ${\mathfrak{F}}$ is a cosheaf on $X$ (with the topology of finite covers) iff for any $U,V\in X$ the following sequence is exact
[…]
where $\alpha=\iota_{U\wedge V,U}-\iota_{U\wedge V,V}$ and $\beta=\iota_{U,U\vee V}+\iota_{V,U\vee V}$.
The following proof is essentially a restatement of the proof of Proposition 1.3 in [16], adapted to the site $L$:
Proof.
We must show that for any $U\in X$ and every covering $U_{1}\vee\ldots\vee U_{n}=U$, the sequence of vector spaces
[…]
is exact, where $\alpha=\sum_{i<j}\iota_{U_{i}\wedge U_{j},U_{i}}-\iota_{U_{i}\wedge U_{j},U_{j}}$ and $\beta=\sum_{i}\iota_{U_{i},U}$. Exactness of the last three terms follows from an easy induction and the associativity of the join operation. For exactness of the first three terms of the sequence, we also proceed by induction: suppose the result holds for all covers of cardinality $n-1$, and let $U_{1},\ldots,U_{n}\in L$ be a cover of $U$. Suppose $(s_{1},\ldots,s_{n})\in\bigoplus_{i=1}^{n}{\mathfrak{F}}(U_{i})$ lies in the kernel of $\beta$. Let $V:=\bigvee_{i=1}^{n-1}U_{i}$. An application of (3) with $U_{n}$ and $V$ shows that $s_{n}=\iota_{U_{n}\cap V,U}(t)$ for some $t\in V\wedge U_{n}=(U_{1}\wedge U_{n})\vee...\vee(U_{n-1}\vee U_{n})$, and by right-exactness of (4) this shows that $s_{n}=\sum_{i=1}^{n-1}\iota_{U_{i}\cap U_{n},U_{n}}(v_{i})$ for some $v_{i}\in{\mathfrak{F}}(U_{i})$. Finally we write
$(s_{1},\ldots,s_{n})=(s_{1}+w_{1},\ldots,s_{n-1}+w_{n-1},0)+(-w_{1},\ldots,-w_{n-1},s_{n})$
where $w_{i}:=\iota_{U_{i}\cap U_{n},U_{n}}(v_{i})$. Both terms above the first term lies in the image of $\alpha$ by the inductive hypothesis, while the second term equals $\sum_{i=1}^{n-1}\iota_{U_{i}\wedge U_{n},U_{n}}(v_{i})-\iota_{U_{i}\wedge U_{n},U_{i}}(v_{i})$, which evidently is also in the image of $\alpha$. ∎
2.2 DGLA attached to a cover
Let $X$ be a distributive pre-ordered set. Let $W\in X$ and ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ be a cover of $W$. Let ${\mathfrak{F}}$ be a pre-cosheaf of vector spaces over $X$. The Čech chain complex $C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})$ is defined by
$C_{n}({\mathfrak{U}},W;{\mathfrak{F}})=\oplus_{i_{0}<\ldots<i_{n}}{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}}),\quad n\geq 0.$
(where some linear order on $I$ has been chosen) with the differential is the Čech differential $\partial:C_{n+1}\rightarrow C_{n}$ given by $\partial=\sum_{j=0}^{n}(-1)^{j}\lambda_{j}$, where $\lambda_{j}$ is the canonical map
${\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}})\to{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\widehat{U_{i_{j}}}\ldots\wedge U_{i_{n}}).$
Lemma 2.2.
If ${\mathfrak{F}}$ is a coflasque cosheaf of vector spaces over $X$, the homology of the Čech complex is 00 for $n>0$ and ${\mathfrak{F}}(W)$ for $n=0$.
Proof.
See [16], Corollary 4.3. Note that [16] uses the term “flabby” instead of “coflasque”. ∎
Let $C^{aug}({\mathfrak{U}},W;{\mathfrak{F}})=\{C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})\rightarrow{\mathfrak{F}}(W)\}$ be the augmented Čech complex.
Proposition 2.1.
Let ${\mathfrak{F}}$ be a local Lie algebra over $X$. Then for any $W\in X$ and any cover ${\mathfrak{U}}$ of $W$ the 1-shifted augmented Čech complex $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ has a natural structure of a non-negatively graded acyclic DGLA.
Proof.
Let ${\mathfrak{U}}$ be a cover of $W$ indexed by $I$. Let $V$ be a vector space with a basis $e_{i}$, $i\in I$ and let $f^{i},i\in I$, be the dual basis of $V^{*}$. The 1-shifted augmented Čech complex of ${\mathfrak{F}}$ with respect to ${\mathfrak{U}}$ is naturally identified with a sub-complex of the DGLA $({\mathfrak{F}}(W)\otimes\Lambda^{\bullet}V,\partial)$, where $\partial$ is contraction with $\sum_{i}f^{i}$. It is easy to check that this sub-complex is closed with respect to the graded Lie bracket thanks to Property I. By Lemma 2.2, the resulting DGLA is acyclic. ∎
Covers of $W\in X$ form a category whose morphisms are refinements. A refinement of a cover ${\mathfrak{U}}=\left\{U_{i}\right\}_{i\in I}$ to a cover ${\mathfrak{V}}=\left\{V_{j}\right\}_{j\in J}$ is a map $\phi:J\rightarrow I$ such that $V_{j}\leq U_{\phi(j)}$. Refinements are composed in an obvious way.
Proposition 2.2.
For a fixed $W\in X$, the map which sends a local Lie algebra ${\mathfrak{F}}$ and a cover ${\mathfrak{U}}$ of $W$ to the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ is functorial in both ${\mathfrak{F}}$ and ${\mathfrak{U}}$.
Proof.
Functoriality in ${\mathfrak{F}}$ is clear. Functoriality in ${\mathfrak{U}}$ is written Čech component-wise for ${\mathsf{p}}^{\mathfrak{V}}\in C_{k}({\mathfrak{V}},W;{\mathfrak{F}})$:
$(\phi_{*}{\mathsf{p}}^{\mathfrak{V}}_{k})_{i_{0},\dots,i_{k}}:=\sum_{j_{0}\in\phi^{-1}(i_{0})}\cdots\sum_{j_{k}\in\phi^{-1}(i_{k})}({\mathsf{p}}^{\mathfrak{V}}_{k})_{j_{0},\dots,j_{k}},$
for $i_{0}<i_{2}<\cdots<i_{k}$. ∎
We will call $C^{aug}_{\bullet+1}(-,W;-)$ the Čech functor. We will need some variants of the Čech functor. First, we can define a graded local Lie algebra over a distributive pre-ordered set $(X,\leq)$ in an obvious manner. The construction of the acyclic DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ works in this case as well, except that it may have components in negative degrees.
Second, we define a pointed DGLA as a DGLA equipped with a distinguished central cycle of degree $-2$ (which we call the curvature). A morphism of pointed DGLAs is a DGLA morphism which preserves the distinguished central cycle. [^5] The category of pointed DGLAs as defined here is a full subcategory of the category of curved DGLAs as defined in [17]. There, for a curved DGLA with curvature ${\mathsf{B}}$, ${\mathsf{B}}$ is not required to be central and the derivation $\partial$ satisfies $\partial^{2}={\rm ad}_{\mathsf{B}}$. We say that a graded local Lie algebra ${\mathfrak{F}}$ over $(X,\leq)$ with a terminal object $T\in X$ is pointed if it is equipped with a distinguished central element ${\mathsf{B}}\in{\mathfrak{F}}(T)$ of degree $-2$. Morphisms in the category of pointed graded local Lie algebras are required to preserve the distinguished element. Then the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}})$ is an acyclic pointed DGLA, the distinguished central cycle being ${\mathsf{B}}\in C^{aug}_{-1}({\mathfrak{U}},T;{\mathfrak{F}})$. Of course, since the DGLA is acyclic, this central cycle is exact. The construction of a pointed DGLA from a pointed graded local Lie algebra and a cover of $T$ is functorial in both arguments.
3 Quantum lattice systems
3.1 Observables and derivations
We use the $\ell^{\infty}$ metric on ${\mathbb{R}}^{n}$, i. e. $d(x,y):=\max_{i=1,\ldots,n}|x_{i}-y_{i}|$. For any $U,V\subset{\mathbb{R}}^{n}$ we write $\operatorname{diam}(U):=\sup_{x,y\in U}d(x,y)$ and $d(U,V):=\inf_{x\in U,y\in V}d(x,y)$. They take values in extended non-negative reals $[0,\infty]$. Thus $\operatorname{diam}(\varnothing)=0$ and $d(U,\varnothing)=\infty$ for any $U\subset{\mathbb{R}}^{n}$. For a nonempty set $U$ and $r\geq 0$ we define $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$ while we set $\varnothing^{r}=\varnothing$.
A quantum lattice system consists of a countable subset $\Lambda\subset{\mathbb{R}}^{n}$ (“the lattice”) and a finite-dimensional complex Hilbert space $V_{j}$ for every $j\in\Lambda$. We make the following assumption on the lattice system [^6] In [11] $\Lambda$ was assumed to be a Delone set, i.e. it was required to be uniformly filling and uniformly discrete. These assumptions were imposed on physical grounds. All the results proved in [11] hold under weaker assumptions adopted in this paper. In [8] $\Lambda$ was taken to be ${\mathbb{Z}}^{n}$ for simplicity. : there is a $C_{\Lambda}>0$ such that the number of points of $\Lambda$ in any hypercube of diameter $d$ is bounded by $C_{\Lambda}(d+1)^{n}$.
For any bounded nonempty $X\subset{\mathbb{R}}^{n}$ let ${\mathscr{A}}(X):=\bigotimes_{j\in X\cap\Lambda}\mathrm{Hom}_{{\mathbb{C}}}(V_{j},V_{j})$. For any $X\subset Y$ there is an inclusion ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ and the algebras ${\mathscr{A}}(X)$ form a direct system with respect to these inclusions. We extend this direct system to include the empty set by setting ${\mathscr{A}}(\varnothing)={\mathbb{C}}$ and letting the inclusion ${\mathscr{A}}(\varnothing)\hookrightarrow{\mathscr{A}}(X)$ take $\alpha\mapsto\alpha\boldsymbol{1}$. Each ${\mathscr{A}}(X)$ is a finite-dimensional $C^{*}$-algebra with the operator norm, and the inclusions ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ preserve this norm. The normed *-algebra of local observables is
[…]
The algebra of quasi-local observables ${\mathscr{A}}$ is the norm-completion of ${\mathscr{A}}_{\ell}$; it is a $C^{*}$-algebra.
For any bounded $X\subset{\mathbb{R}}^{n}$, define the normalized trace ${\overline{\rm tr}}:{\mathscr{A}}(X)\to{\mathbb{C}}$ as ${\overline{\rm tr}}({\mathcal{A}})=\operatorname{tr}({\mathcal{A}})/{\sqrt{\dim({\mathscr{A}}(X))}}$. For any bounded $X\subset Y$ the partial trace ${\overline{\rm tr}}_{X^{c}}:{\mathscr{A}}(Y)\to{\mathscr{A}}(X)$ is uniquely specified by the condition ${\overline{\rm tr}}_{X^{c}}({\mathcal{A}}\otimes{\mathcal{B}})={\overline{\rm tr}}({\mathcal{A}}){\mathcal{B}}$ for any ${\mathcal{A}}\in{\mathscr{A}}(Y\backslash X)$ and ${\mathcal{B}}\in{\mathscr{A}}(X)$. Besides forming a direct system with respect to inclusions, the spaces ${\mathscr{A}}(X)$ are also an inverse system with respect to the partial trace. ${\overline{\rm tr}}$ extends to a normalized positive linear functional on ${\mathscr{A}}$, i.e. a state. We say that ${\mathcal{A}}\in{\mathscr{A}}$ is traceless if ${\overline{\rm tr}}({\mathcal{A}})=0$. The space of traceless anti-hermitian elements of ${\mathscr{A}}(X)$ will be denoted ${{\mathfrak{d}}_{l}}(X)$. ${{\mathfrak{d}}_{l}}(X)$ is a real Lie algebra with respect to the commutator. The Lie algebras ${{\mathfrak{d}}_{l}}(X)$ form a direct system over the directed set of bounded subsets of ${\mathbb{R}}^{n}$, and its limit will be denoted ${{\mathfrak{d}}_{l}}$. Equivalently, ${{\mathfrak{d}}_{l}}$ is the Lie algebra of traceless anti-hermitian elements of ${\mathscr{A}}_{\ell}$. Note that ${\mathscr{A}}_{\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{l}}\otimes{\mathbb{C}})$.
Definition 3.1.
A brick in ${\mathbb{R}}^{n}$ is a non-empty subset of the form
$Y=\{(x_{1},\ldots,x_{n})\ |\ \ell_{i}\leq x_{i}\leq m_{i},i=1,\ldots,n\}$ (5)
where $(k_{1},\ldots,k_{n})$, $(\ell_{1},\ldots,\ell_{n})$, and $(m_{1},\ldots,m_{n})$ are $n$-tuples of integers. We write ${\mathbb{B}}_{n}$ for the set of all bricks in ${\mathbb{R}}^{n}$.
The set of bricks exhausts the collection of bounded subsets of ${\mathbb{R}}^{n}$ in the sense that any bounded subset is contained in a brick. In addition, the set of bricks satisfies the following regularity property:
Lemma 3.1.
For any $j\in{\mathbb{R}}^{n}$ we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\frac{\pi^{4}4^{n}(n+1)^{2}}{36}$
Proof.
Any pair of points $x,y\in{\mathbb{Z}}^{n}$ specifies a brick with $x$ and $y$ on opposing corners, and any brick can be specified this way (not uniquely). With $X$ the brick corresponding to $x$ and $y$ it is easy to see that $\max(d(x,j),d(y,j))\leq\operatorname{diam}(X)+d(X,j)$, and so $(1+d(x,j))(1+d(y,j))\leq(1+\operatorname{diam}(X)+d(X,j))^{2}$. Thus we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\sum_{\begin{subarray}{c}x,y\in{\mathbb{Z}}^{n}\end{subarray}}(1+d(x,j))^{-n-1}(1+d(y,j))^{-n-1}$
$=\left(\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,j))^{-n-1}\right)^{2},$
and it remains only to bound the above sum. Let $f(k):=(1+k)^{-n-1}$ and $g(k):=\#({\mathbb{Z}}^{n}\cap B_{k}(j))\leq(1+2k)^{n}$. Using summation by parts we have
$\sum_{j\in\Lambda}(1+d(x,j))^{-n-1}$ $\leq\sum_{k\geq 0}f(k)(g(k+1)-g(k))$
$=\lim_{k\to\infty}f(k)g(k)-\sum_{k\geq 0}g(k)(f(k+1)-f(k)).$
It is easy to check that $f(k)g(k)\to 0$ and that $-(f(k+1)-f(k))\leq(n+1)(1+k)^{-n-2}$, and so
$\sum_{j\in\Lambda}(1+d(x,j))^{-n-1}$ $\leq 2^{n}(n+1)\sum_{k\geq 0}(1+k)^{-2}$
$\leq\frac{\pi^{2}2^{n}(n+1)}{6}$
which proves the Lemma. ∎
For any brick $Y$ we define the following subspace of ${{\mathfrak{d}}_{l}}(Y)$:
${{\mathfrak{d}}_{l}}^{Y}:=\{{\mathcal{A}}\in{{\mathfrak{d}}_{l}}(Y)\ |\ {\overline{\rm tr}}_{X^{c}}({\mathcal{A}})=0\text{ for any brick }X\subsetneq Y\}.$
Each ${{\mathfrak{d}}_{l}}(Y)$ decomposes as a direct sum ${{\mathfrak{d}}_{l}}(Y)=\bigoplus_{X\subseteq Y}{{\mathfrak{d}}_{l}}^{X}$ over bricks contained in $Y$, and for any brick $X\subseteq Y$ the partial trace ${\overline{\rm tr}}_{Y\backslash X^{c}}$ is the projection onto $\bigoplus_{Z\subseteq X}{{\mathfrak{d}}_{l}}^{Z}\subseteq{{\mathfrak{d}}_{l}}(Y)$. Intuitively, ${{\mathfrak{d}}_{l}}^{Y}$ consists of elements of ${{\mathfrak{d}}_{l}}(Y)$ which are not localized on any brick properly contained in $Y$.
Derivations of ${\mathscr{A}}$ which appear in the physical context are typically only densely defined and have the form
${\mathsf{F}}:{\mathcal{A}}\mapsto\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}],$
where ${\mathsf{F}}^{Y}\in{{\mathfrak{d}}_{l}}^{Y}$. We are now going to define for every $U\subset{\mathbb{R}}^{n}$ a real Lie algebra ${\mathfrak{D}}_{al}(U)$ which consists of derivations approximately localized on $U$ and such that all ${\mathfrak{D}}_{al}(U)$ have a common dense domain. Moreover, they form a pre-cosheaf of Lie algebras over a certain category of subsets of ${\mathbb{R}}^{n}$.
Definition 3.2.
For any element ${\mathsf{F}}=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$ of $\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ and every $U\subset{\mathbb{R}}^{n}$ we let
$\|{\mathsf{F}}\|_{U,k}:=\sup_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|(1+\operatorname{diam}(Y)+d(U,Y))^{k}$ (6)
and define ${\mathfrak{D}}_{al}(U)\subset\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ as the set of elements ${\mathsf{F}}$ with $\|{\mathsf{F}}\|_{U,k}<\infty$ for all $k\geq 0$.
If $U$ is empty, then for $k>0$ $\|{\mathsf{F}}\|_{U,k}<\infty$ if and only if ${\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. Thus ${\mathfrak{D}}_{al}(\emptyset)=0$. We also denote ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})={\mathfrak{D}}_{al}$.
It is easy to see that (6) is a norm on ${\mathfrak{D}}_{al}(U)$ for each $k\geq 0$. We endow ${\mathfrak{D}}_{al}(U)$ with the locally convex topology given by the norms (6) ranging over all $k\geq 0$. Recall that a topological vector space is called a Fréchet space if it is Hausdorff, and if its topology can be generated by a countable family of seminorms with respect to which it is complete.
Proposition 3.1.
${\mathfrak{D}}_{al}(U)$ is a Fréchet space.
Proof.
The Hausdorff property follows from the fact that if $\|{\mathsf{F}}\|_{U,k}=0$ for any $k\geq 0$ then ${\mathsf{F}}=0$. To show completeness, suppose $\{{\mathsf{F}}_{m}\}_{m\in\mathbb{N}}\subset{\mathfrak{D}}_{al}(U)$ is Cauchy, i.e. that for any $k\geq 0$ and any $\epsilon>0$ there is an $N\in\mathbb{N}$ such that $m,m^{\prime}\geq N\implies\|{\mathsf{F}}_{m}-{\mathsf{F}}_{m^{\prime}}\|_{U,k}<\epsilon$. For any fixed $Y\in{\mathbb{B}}_{n}$ this implies that $\{{\mathsf{F}}_{n}^{Y}\}$ is Cauchy in ${{\mathfrak{d}}_{l}}^{Y}$ (with the operator norm) and thus converges to a limit ${\mathsf{F}}^{Y}$. Let ${\mathsf{F}}:=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$.
Fix $k\in\mathbb{N}$ and $\epsilon>0$. For every $\ell=0,1,2,...$, choose $N_{\ell}\in\mathbb{N}$ so that $m,m^{\prime}\geq N_{\ell}\implies\|{\mathsf{F}}_{m}-{\mathsf{F}}_{m^{\prime}}\|_{U,k}<2^{-\ell-1}\epsilon$. For any $Y\in{\mathbb{B}}_{n}$, any $m\geq N_{1}$, and any $M>1$, we have
$\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}_{N_{1}}^{Y}\|+\sum_{i=1}^{M-1}\|{\mathsf{F}}_{N_{i}}^{Y}-{\mathsf{F}}_{N_{i+1}}^{Y}\|+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|$
$\leq\epsilon(1+\operatorname{diam}(Y)+d(U,Y))^{-k}+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|.$
Taking $M\to\infty$ shows that $\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}^{Y}\|(1+\operatorname{diam}(Y)+d(U,Y))^{k}<\epsilon$. Since $Y$ was arbitrary, we have $\|{\mathsf{F}}_{m}-{\mathsf{F}}\|_{U,k}<\epsilon$. Since $k$ was arbitrary, $\{{\mathsf{F}}_{m}\}$ converges to ${\mathsf{F}}$ in the topology of ${\mathfrak{D}}_{al}(U)$. ∎
Recall that for a nonempty $U\subset{\mathbb{R}}^{n}$ we write $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$. The norms $\|\cdot\|_{U,k}$ obey the following dominance relation.
Lemma 3.2.
Let $U,V$ be subsets of the lattice and suppose that $U\subseteq V^{r}$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ we have
$\|{\mathsf{F}}\|_{V,k}\leq(r+1)^{k}\|{\mathsf{F}}\|_{U,k}.$ (7)
In particular, ${\mathfrak{D}}_{al}(U)\subseteq{\mathfrak{D}}_{al}(V)$ and the inclusion is continuous.
Proof.
Let $Y$ be any subset of the lattice. Then (7) follows from
$1+\operatorname{diam}(Y)+d(Y,V)$ $\leq 1+\operatorname{diam}(Y)+d(Y,V^{r})+r$
$\leq 1+\operatorname{diam}(Y)+d(Y,U)+r$
$\leq(r+1)(1+\operatorname{diam}(Y)+d(Y,U)),$
where in the second line we used the triangle inequality and in the third we used $U\subseteq V^{r}$. ∎
The above Lemma shows that the space ${\mathfrak{D}}_{al}(U)$ only depends on the asymptotic geometry of the region $U$ in the following sense: if $U\subseteq V^{r}$ and $V\subseteq U^{r}$ for some $r\geq 0$ then ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(V)$ (as subsets of $\prod_{Y}{{\mathfrak{d}}_{l}}^{Y})$ and are isomorphic as Fréchet spaces. In particular, for any non-empty bounded $U\subset{\mathbb{R}}^{n}$, the space ${\mathfrak{D}}_{al}(U)$ coincides with ${\mathfrak{D}}_{al}(\{0\})$.
To relate the spaces ${\mathfrak{D}}_{al}(U)$ to the traditional $C^{*}$-algebraic picture we prove the following:
Proposition 3.2.
Suppose $U\subset{\mathbb{R}}^{n}$ is non-empty and bounded and let ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$. Then the sum $\sum_{X\in{\mathbb{B}}_{n}}{\mathsf{F}}^{X}$ is absolutely convergent and defines a continuous dense embedding of ${\mathfrak{D}}_{al}(U)$ into the subspace of traceless anti-hermitian elements of the algebra ${\mathscr{A}}$.
Proof.
We can assume without loss of generality that $U=\{0\}$. We have
$\sum_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}\|_{\{0\},2n+2}\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,0))^{-2n-2},$
and by Lemma 3.1 the above sum is finite. This shows that the map ${\mathfrak{D}}_{al}(\{0\})\to{\mathscr{A}}$ is well-defined and continuous. Its image is dense in the space of traceless anti-hermitian elements of ${\mathscr{A}}$ because it contains the dense subspace ${{\mathfrak{d}}_{l}}$. Finally, to show that it is injective, writing ${\mathcal{A}}:=\sum_{Y\in{\mathbb{B}}_{n}}{\mathsf{F}}^{Y}$ we have
${\mathsf{F}}^{Y}={\overline{\rm tr}}_{Y^{c}}({\mathcal{A}})-\sum_{X\subsetneq Y}{\overline{\rm tr}}_{X^{c}}({\mathcal{A}})$
and so ${\mathcal{A}}=0\implies{\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. ∎
Definition 3.3.
We let ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}$ be the image of ${\mathfrak{D}}_{al}(\{0\})$ under the embedding of Prop. 3.2, with the Fréchet topology of ${\mathfrak{D}}_{al}(\{0\})$.
In [11], the algebra ${\mathscr{A}}_{a\ell}$ of almost-local operators was defined as a subspace of ${\mathscr{A}}$ where a countable family of norms similar to (6) takes finite values. Here we equivalently define ${\mathscr{A}}_{a\ell}$ as the set of elements of ${\mathscr{A}}$ whose traceless hermitian and anti-hermitian parts live in ${{\mathfrak{d}}_{al}}$, topologized as ${\mathscr{A}}_{a\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{al}}\otimes{\mathbb{C}})$.
Proposition 3.3.
${\mathscr{A}}_{a\ell}$ is a dense sub-algebra of ${\mathscr{A}}$.
Proof.
${\mathscr{A}}_{a\ell}$ contains all local observables and these are dense in ${\mathscr{A}}$. The fact that ${\mathscr{A}}_{a\ell}$ is closed under multiplication is proven in [11]. ∎
In view of Prop. 3.2 it is natural to make the following definition.
Definition 3.4.
An element ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ is inner iff it is contained in ${\mathfrak{D}}_{al}(U)$ for some bounded $U$.
For any two bounded sets $U,V$, a local observable is strictly localized on both $U$ and $V$ iff it is localized on their intersection, ie. ${\mathscr{A}}(U)\cap{\mathscr{A}}(V)={\mathscr{A}}(U\cap V)$. An analogous relation for the spaces ${\mathfrak{D}}_{al}(U)$ does not hold in general [^7] Indeed, for any two bounded $U,V$ the spaces ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}(V)$ coincide and are nontrivial but if $U$ and $V$ are disjoint then ${\mathfrak{D}}_{al}(U\cap V)=0$., but it does hold if we assume that $U$ and $V$ satisfy the following transversality condition:
Definition 3.5.
Let $U,V\subseteq{\mathbb{R}}^{n}$ and $C>0$. We say $U$ and $V$ are $C$-transverse if
$d(x,U\cap V)\leq C\max(d(x,U),d(x,V))$
for all $x\in{\mathbb{R}}^{n}$.
We will say $U$ and $V$ are transverse if they are $C$-transverse for some $C>0$.
Proposition 3.4.
If $U,V\subseteq{\mathbb{R}}^{n}$ are $C$-transverse then
$\max(\|{\mathsf{F}}\|_{U,k},\|{\mathsf{F}}\|_{V,k})\leq\|{\mathsf{F}}\|_{U\cap V,k}\leq(C+1)^{k}\max(\|{\mathsf{F}}\|_{U,k},\|{\mathsf{F}}\|_{V,k})$ (8)
for all $k>0$. In particular, ${\mathfrak{D}}_{al}(U\cap V)$ is a topological pullback: it is the set ${\mathfrak{D}}_{al}(U)\cap{\mathfrak{D}}_{al}(V)$ with the topology of simultaneous convergence in ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}(V)$.
Proof.
The first inequality is true even without assuming transversality – it follows from Lemma 3.2. For the second, let $Z\in{\mathbb{B}}_{n}$ and choose $x^{*},y^{*}\in Z$ so that $d(x^{*},U)=d(Z,U)$ and $d(y^{*},V)=d(Z,V)$. Then we have
$d(Z,U\cap V)$ $=\inf_{z\in Z}d(z,U\cap V)$
$\leq C\inf_{z\in Z}\max(d(z,U),d(z,V))$
$\leq C\inf_{z\in Z}\max(d(z,x^{*})+d(x^{*},U),d(z,y^{*})+d(y^{*},V))$
$\leq C(\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V)),$
and thus
$1+\operatorname{diam}(Z)+d(Z,U\cap V)\leq$
$\hskip 28.45274pt\leq(C+1)(1+\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V))),$ (9)
which proves (8). ∎
The next proposition relates ${\mathfrak{D}}_{al}(U\cup V)$ with ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}(V)$ for any $U,V\subseteq{\mathbb{R}}^{n}$.
Proposition 3.5.
For any $U,V\subset{\mathbb{R}}^{n}$ consider the sequence of vector spaces
[…]
where $\alpha({\mathsf{F}})=({\mathsf{F}},-{\mathsf{F}})$ and $\beta({\mathsf{F}},{\mathsf{G}})={\mathsf{F}}+{\mathsf{G}}$.
  1. i)
    The sequence (10) admits a right splitting, i.e. a map $\gamma:{\mathfrak{D}}_{al}(U\cup V)\to{\mathfrak{D}}_{al}(U)\oplus{\mathfrak{D}}_{al}(V)$ with $\beta\circ\gamma=\operatorname{id}$. In particular, it is exact on the right.
  2. ii)
    If $U$ and $V$ are transverse, then (10) is exact on the left.
Proof.
$i)$. For any $Y\in{\mathbb{B}}_{n}$, define
$\chi(Y):=\left\{\begin{array}[]{cc}1&\text{ if }d(Y,U)<d(Y,V)\\ 1/2&\text{ if }d(Y,U)=d(Y,V)\\ 0&\text{ if }d(Y,U)>d(Y,V)\end{array}\right..$
For any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U\cup V)$ define $\gamma_{1}({\mathsf{F}}):=\sum_{Y\in{\mathbb{B}}_{n}}\chi(Y){\mathsf{F}}^{Y}$, and $\gamma_{2}({\mathsf{F}})={\mathsf{F}}-\gamma_{1}({\mathsf{F}})$. Then it is not hard to show that $\|\gamma_{1}({\mathsf{F}})\|_{U,k}\leq\|{\mathsf{F}}\|_{U\cup V,k}$ and $\|\gamma_{2}({\mathsf{F}})\|_{V,k}\leq\|{\mathsf{F}}\|_{U\cup V,k}$ and that $\gamma=(\gamma_{1},\gamma_{2})$ is a right splitting of (10).
$ii)$. Follows immediately from Proposition 3.4. ∎
Next we will show how to endow ${\mathfrak{D}}_{al}(U)$ with the Lie algebra structure.
Proposition 3.6.
Let $U,V\subset{\mathbb{R}}^{n}$. For any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ and ${\mathsf{G}}\in{\mathfrak{D}}_{al}(V)$ the sum
$[{\mathsf{F}},{\mathsf{G}}]^{Z}:=\sum_{X,Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}$ (11)
is absolutely convergent for every $Z\in{\mathbb{B}}_{n}$. The resulting bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] satisfies the Jacobi identity and
$\|[{\mathsf{F}},{\mathsf{G}}]\|_{U,k}$ $\leq C3^{k}\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{V,k+4n+4}$ (12)
for some constant $C>0$ that depends only on $n$.
To prove Proposition 3.6 we will need several lemmas. For any $X,Y\in{\mathbb{B}}_{n}$, define [^8] This is well-defined since the intersection of an arbitrary number of bricks is either empty or a brick, so $X\vee Y$ is the intersection of all bricks containing $X$ and $Y$. the join $X\vee Y\in{\mathbb{B}}_{n}$ as the smallest brick that contains $X$ and $Y$.
Lemma 3.3.
For any $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ we have
$\operatorname{diam}(X\vee Y)\leq\operatorname{diam}(X)+\operatorname{diam}(Y)$ (13)
and for any $z\in X\vee Y$ we have
$d(z,X)\leq\operatorname{diam}(Y).$
Proof.
Let $\pi_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ be the projection onto the $i$th coordinate. The following identites hold for any bricks $X,Y\in{\mathbb{B}}_{n}$:
$\pi_{i}(X\vee Y)$ $=\pi_{i}(X)\vee\pi_{i}(Y),$
$\operatorname{diam}(X)$ $=\max_{i=1,\ldots n}\operatorname{diam}(\pi_{i}(X)),$
$d(X,Y)$ $=\max_{i=1,\ldots n}d(\pi_{i}(X),\pi_{i}(Y)).$
When $n=1$ the results are clear. When $n>1$ they follow from the $n=1$ case via the above identities. ∎
Lemma 3.4.
Let $X,Y,Z\in{\mathbb{B}}_{n}$ and let ${\mathsf{F}}\in{{\mathfrak{d}}_{l}}(X)$ and ${\mathsf{G}}\in{{\mathfrak{d}}_{l}}(Y)$. Then $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}=0$ unless $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$.
Proof.
The requirement that $X\cap Y\neq\varnothing$ is clear, since ${\mathsf{F}}^{X}$ and ${\mathsf{G}}^{Y}$ would commute otherwise. Suppose $Z\nsubseteq X\vee Y$. Then $Z^{\prime}:=(X\vee Y)\cap Z$ is a brick that is strictly contained in $Z$, so $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\in{{\mathfrak{d}}_{l}}^{Z}\cap{{\mathfrak{d}}_{l}}(Z^{\prime})=\{0\}.$
We make the following definitions for the next lemma. For any $U\subset{\mathbb{R}}^{n}$ and any brick $X$ write $\tilde{d}(X,U):=1+\operatorname{diam}(X)+d(X,U)$ and $\tilde{d}(X):=1+\operatorname{diam}(X)$.
Lemma 3.5.
For any $U\subset{\mathbb{R}}^{n}$, $Z\in{\mathbb{B}}_{n}$, and $k\geq 0$ we have
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq\frac{\pi^{8}16^{n}(n+1)^{4}3^{k}}{1296}\tilde{d}(Z,U)^{-k}.$
Proof.
Let $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ and let $Z\subset X\vee Y$ be a brick. Pick an arbitrary point $w\in X\cap Y$. We have
$d(Z,U)$ $\leq d(Z,w)+d(w,U)$
$\leq d(Z,w)+\operatorname{diam}(X)+d(X,U)$
$\leq\operatorname{diam}(X\vee Y)+\operatorname{diam}(X)+d(X,U)$
$\leq 2\operatorname{diam}(X)+\operatorname{diam}(Y)+d(X,U),$
and so, since by Lemma 3.3 $\operatorname{diam}(Z)\leq\operatorname{diam}(X\vee Y)\leq\operatorname{diam}(X)+\operatorname{diam}(Y)$, we have
$\tilde{d}(Z,U)$ $\leq 1+3\operatorname{diam}(X)+2\operatorname{diam}(Y)+d(X,U)$
$\leq 3(1+\operatorname{diam}(X)+d(X,U)+\operatorname{diam}(Y))$
$\leq 3(1+\operatorname{diam}(X)+d(X,U))(1+\operatorname{diam}(Y))$
$=3\tilde{d}(X,U)\tilde{d}(Y).$ (14)
It follows that for any $k>0$ we have
$\tilde{d}(Z,U)^{k}\tilde{d}(X,U)^{-k}\tilde{d}(Y)^{-k}$ $\leq 3^{k}$
and so
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq$
$\hskip 28.45274pt\leq 3^{k}\tilde{d}(Z,U)^{-k}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ (15)
It remains to bound the sum on the right-hand side. Fix an arbitrary $z\in Z$ and let $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ and $Z\subset X\vee Y$. Then by the second statement in Lemma 3.3 and the inequality $1+a+b\leq(1+a)(1+b)$ for $a,b\geq 0$ we have
$(1+d(z,X)+\operatorname{diam}(X))(1+d(z,Y)+\operatorname{diam}(Y))\leq$
$\hskip 42.67912pt\leq(1+\operatorname{diam}(X))^{2}(1+\operatorname{diam}(Y))^{2}.$ (16)
Using (16) we can bound the sum (15) as follows:
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}\leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}(1+\operatorname{diam}(X))^{-4n-4}(1+\operatorname{diam}(Y))^{-4n-4}\\ \leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}\\ \leq\left(\sum_{\begin{subarray}{c}X\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}\right)^{2}\leq\frac{\pi^{8}16^{n}(n+1)^{4}}{1296}.$ (17)
Now we are ready to prove Proposition 3.6.
Proof of Proposition 3.6.
Notice that by Lemma 3.2 we have $\|{\mathsf{G}}\|_{V,k+2n+2}\leq\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+2n+2}$ so without loss of generality we set $V={\mathbb{R}}^{n}$.
Let $Z\in{\mathbb{B}}_{n}$. By Lemmas 3.4 and 3.5 we have
$\sum_{X,Y\in{\mathbb{B}}_{n}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$ $=\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$
$\leq 2\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}$
$\leq C3^{k}\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\tilde{d}(Z,U)^{-k},$
with $C=\frac{\pi^{8}16^{n}(n+1)^{4}}{648}$. This proves that (11) is absolutely convergent and establishes the bound (12).
Next, let us prove the Jacobi identity. Let ${\mathsf{F}},{\mathsf{G}},{\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$. We have
$[{\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}$ $=\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\left[{\mathsf{F}}^{X},\left[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}\right]^{Y}\right]^{W}$ (18)
To show that this sum is absolutely convergent, we bound
$\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\left\|\left[{\mathsf{F}}^{X},\left[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}\right]^{Y}\right]^{W}\right\|$
$\leq$ $4\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\sum_{\begin{subarray}{c}X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Y\subset X^{\prime}\vee Y^{\prime}\end{subarray}}\|{\mathsf{G}}^{X^{\prime}}\|\|{\mathsf{H}}^{Y^{\prime}}\|$
$\leq$ $C4\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\tilde{d}(Y)^{-4n-4}$
$\leq$ $C4\|{\mathsf{F}}\|_{{\mathbb{R}}^{n},4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\tilde{d}(X)^{-4n-4}\tilde{d}(Y)^{-4n-4}$
$<$ $\infty$.∞ .
Here we used Lemma 3.5 in the second and fourth lines and $C$ is a constant depending only on $n$. Thus the sum (18) is absolutely convergent. In particular, using the fact that $\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}=[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]$ we have the following absolutely convergent expression
$[{\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}=\sum_{X,Y,Z\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},[{\mathsf{G}}^{Y},{\mathsf{H}}^{Z}]]^{W}.$
It is then easy to check that the Jacobi identity for the sum follows from the Jacobi identity for each term. ∎
From Propositions 3.4 and 3.6 we immediately get
Corollary 3.1.
Suppose $U,V\in{\mathbb{R}}^{n}$.
  1. i)
    $({\mathsf{F}},{\mathsf{G}})\mapsto[{\mathsf{F}},{\mathsf{G}}]$ is a jointly continuous bilinear map from ${\mathfrak{D}}_{al}(U)\times{\mathfrak{D}}_{al}(V)$ to ${\mathfrak{D}}_{al}(U)\cap{\mathfrak{D}}_{al}(V)$.
  2. ii)
    If $U$ and $V$ are transverse, then this is jointly continuous bilinear map from ${\mathfrak{D}}_{al}(U)\times{\mathfrak{D}}_{al}(V)$ to ${\mathfrak{D}}_{al}(U\cap V)$.
In particular since $\{0\}$ and ${\mathbb{R}}^{n}$ are transverse, ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ acts continuously on ${\mathfrak{D}}_{al}(\{0\})$ and this action is easily seen to extend to a continuous action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on the space ${\mathscr{A}}_{a\ell}$. We denote the action of ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ by ${\mathcal{A}}\mapsto{\mathsf{F}}({\mathcal{A}})$. By Prop. 3.6, it is given by
${\mathsf{F}}({\mathcal{A}})=\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}].$ (19)
It is not hard to check that for any $X\in{\mathbb{B}}_{n}$ and any ${\mathcal{A}}\in{\mathscr{A}}(X)$ we have ${\overline{\rm tr}}_{X^{c}}\left({\mathsf{F}}({\mathcal{A}})\right)=[{\mathsf{F}}_{X},{\mathcal{A}}]$ and so the action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathscr{A}}_{a\ell}$ is faithful. Thus, elements of ${\mathfrak{D}}_{al}(U)$ for any $U\subset{\mathbb{R}}^{n}$ may be identified with a subset of the Fréchet-continuous derivations of ${\mathscr{A}}_{a\ell}$. By Proposition 3.2, the Fréchet-Lie algebra ${\mathfrak{D}}_{al}(\{0\})$ is identified with the Fréchet-Lie algebra ${{\mathfrak{d}}_{al}}$ of traceless anti-hermitian elements of ${\mathscr{A}}_{a\ell}$ acting by inner derivations.
3.2 Automorphisms
In this section we recall certain automorphisms obtained by exponentiating elements of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ following [11]. One can develop the theory of such automorphisms that are almost-localized on regions in ${\mathbb{R}}^{n}$ in a similar spirit to the above, but since we do not have much occasion to use them in this work, we opt instead for a more minimal development. Let ${\mathsf{F}}:{\mathbb{R}}\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ be a smooth map. It is shown in [11] that for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ the differential equation
$\frac{d}{dt}{\mathcal{A}}(t)={\mathsf{F}}(t)({\mathcal{A}}(t))$
with the initial condition ${\mathcal{A}}(0)={\mathcal{A}}$ has a unique solution ${\mathcal{A}}(t)\in{\mathscr{A}}_{a\ell}$ for all $t\in{\mathbb{R}}$. Denote by $\alpha^{\mathsf{F}}_{t}$ the map taking ${\mathcal{A}}$ to ${\mathcal{A}}(t)$. It is a continuous automorphism of the Lie algebra ${\mathscr{A}}_{a\ell}$ that preserves the *-operation.
Definition 3.6.
We call any automorphism of the form $\alpha^{\mathsf{F}}_{t}$ for some smooth map ${\mathsf{F}}:{\mathbb{R}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ a locally-generated automorphism, or LGA for short.
It is shown in [11] that every LGA extends to a continuous *-automorphism of the quasilocal algebra ${\mathscr{A}}$, and that the set of LGAs forms a group under composition.
3.3 States
By a state $\psi$ we will mean a state on the quasilocal algebra ${\mathscr{A}}$. If ${\mathsf{F}}$ is an inner derivation (see Def. 3.4), we define its $\psi$-average as the evaluation of $\psi$ on the corresponding element of ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}_{a\ell}$. The group of LGAs acts on states by pre-composition, which we denote $\psi^{\alpha}:=\psi\circ\alpha$. We say an LGA $\alpha$ preserves a state $\psi$ if $\psi^{\alpha}=\psi$. We say an element ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ preserves $\psi$ if $\psi({\mathsf{F}}({\mathcal{A}}))=0$ for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$, which is equivalent to the one-parameter group of automorphisms $t\mapsto\alpha_{t}^{\mathsf{F}}$ corresponding to a constant map $t\mapsto{\mathsf{F}}$ preserving $\psi$.
Definition 3.7.
For any $U\subset{\mathbb{R}}^{n}$ define ${\mathfrak{D}}_{al}^{\psi}(U)$ as the set of all elements of ${\mathfrak{D}}_{al}(U)$ that preserve $\psi$.
It is easy to check that ${\mathfrak{D}}_{al}^{\psi}(U)$ is a closed subset of ${\mathfrak{D}}_{al}(U)$, and that if ${\mathsf{F}}$ and ${\mathsf{G}}$ preserve $\psi$, then $[{\mathsf{F}},{\mathsf{G}}]$ preserves $\psi$. Thus the analog of Propositions 3.4 and 3.6 and Corollary 3.1 hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$. Proposition 3.5 on the other hand, does not hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$ for a general state $\psi$. To circumvent this, we will restrict to gapped states, where quasiadiabatic evolution [18, 19, 20] can be used to prove the analog of Proposition 3.5.
Definition 3.8.
A state $\psi$ is gapped if there exists ${\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ and $\Delta>0$ such that for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ one has
$-i\psi({\mathcal{A}}^{*}{\mathsf{H}}({\mathcal{A}}))\geq\Delta\left(\psi({\mathcal{A}}^{*}{\mathcal{A}})-\psi({\mathcal{A}}^{*})\psi({\mathcal{A}})\right).$ (20)
Remark 3.1.
The meaning of this condition becomes more transparent if one recalls that any ${\mathsf{H}}\in{\mathfrak{D}}_{al}$ is a generator of a one-parameter group of *-automorphisms of ${\mathscr{A}}$ [11]. The condition (20) implies that $\psi$ is invariant under this one-parameter group of automorphisms [21], and that the corresponding one-parameter group of unitaries in the GNS representation of ${\mathscr{A}}$ has a generator whose spectrum in the orthogonal complement to the GNS vacuum vector is contained in $[\Delta,+\infty)$. The condition (20) also implies that $\psi$ is pure [22].
Remark 3.2.
If $\psi$ is a gapped state of ${\mathscr{A}}$ and $\alpha$ is an LGA, then $\psi^{\alpha}$ is also a gapped state. In [11] it was proposed to define a gapped phase as an orbit of gapped state under the action of the group of LGAs.
In Appendix A we prove the following.
Proposition 3.7.
Suppose $\psi$ is gapped, the corresponding Hamiltonian is ${\mathsf{H}}$. Then there are linear functions
$\mathcal{J}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}^{\psi}({\mathbb{R}}^{n})$
$\mathcal{K}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$
such that
  1. i)
    If ${\mathsf{F}}$ preserves $\psi$ then $\mathcal{K}({\mathsf{F}})$ preserves $\psi$.
  2. ii)
    For every $k>0$, $U\subset{\mathbb{R}}^{n}$, and ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ we have
    $\|\mathcal{J}({\mathsf{F}})\|_{U,k}$ $\leq C_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$
    $\|\mathcal{K}({\mathsf{F}})\|_{U,k}$ $\leq C^{\prime}_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$
    for some constants $C_{k},C^{\prime}_{k}$ depending only on $k,n,{\mathsf{H}}$, and $\Delta$.
  3. iii)
    For every ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ we have
    ${\mathsf{F}}=\mathcal{J}({\mathsf{F}})-\mathcal{K}([{\mathsf{H}},{\mathsf{F}}]).$
Using the above Lemma we will prove the analog of Proposition 3.5 for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$.
Proposition 3.8.
For any $U,V\subset{\mathbb{R}}^{n}$ consider the sequence
[…]
where $\alpha({\mathcal{A}})=({\mathcal{A}},-{\mathcal{A}})$ and $\beta({\mathcal{A}},{\mathcal{B}})={\mathcal{A}}+{\mathcal{B}}$.
  1. i)
    If $\psi$ is gapped then the sequence (10) admits a right splitting, and in particular it is exact on the right
  2. ii)
    If $U$ and $V$ are transverse, then (10) is exact on the left.
To prove Proposition 3.8 we will need the following geometric result
Lemma 3.6.
Let $U,V\subset{\mathbb{R}}^{n}$ and define $U^{\prime}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\}$. Then $U^{\prime}$ and $U\cup V$ are transverse and their intersection is $U$.
Proof.
It is easy to check that $U^{\prime}\cap(U\cup V)=U$. To prove transversality we will show
$d(x,U)\leq 4\max(d(x,U^{\prime}),d(x,U\cup V))$ (22)
for every $x\in{\mathbb{R}}^{n}$. Suppose first that $d(x,U)\leq 2d(x,V)$. Then
$d(x,U)$ $\leq 2\min(d(x,U),d(x,V))$
$=2d(x,U\cup V)$
which implies (22). Suppose instead that $d(x,U)>2d(x,V)$, and let $y\in U^{\prime}$ satisfy $d(x,y)=d(x,U^{\prime})$. Notice $x\notin U^{\prime}$ and so $y$ lies in the boundary of $U^{\prime}$, which implies $d(y,U)=d(y,V)$. Thus we have
$d(x,U)$ $\leq d(x,y)+d(y,U)$
$=d(x,y)+d(y,V)$
$\leq 2d(x,y)+d(x,V)$
where in the first and third lines we used the triangle inequality. Using $d(x,y)=d(x,U^{\prime})$ and $d(x,V)<d(x,U)/2$, this gives $d(x,U)<4d(x,U^{\prime})$, which implies (22). ∎
Proof of Proposition 3.8.
The proof of Proposition 3.5 goes through unmodified except for the definition of $\gamma$, which needs to be changed to ensure that the image of $\gamma$ consists of derivations that preserve $\psi$. Suppose $U,V\subseteq{\mathbb{R}}^{n}$ and ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U\cup V)$. Define
$U^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\},$
$V^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,V)\leq d(x,U)\}.$
Let $\gamma^{U,V}$ (resp. $\gamma^{U^{\prime},V^{\prime}}$) be the splitting from Proposition 3.5 with the sets $U$ and $V$ (resp. $U^{\prime}$ and $V^{\prime}$). Define $\tilde{\gamma}=(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$ as
$\tilde{\gamma}_{i}({\mathsf{F}})$ $:=\mathcal{J}(\gamma_{i}^{U,V}({\mathsf{F}}))-\mathcal{K}([\mathcal{J}(\gamma_{i}^{U^{\prime},V^{\prime}}({\mathsf{H}})),{\mathsf{F}}])$
for $i=1,2$. Using Prop. 3.7 and Lemma 3.6 and the fact that the commutator of two derivations that preserve $\psi$ preserves $\psi$, one checks that $\tilde{\gamma}$ takes ${\mathfrak{D}}_{al}^{\psi}(U\cup V)$ to ${\mathfrak{D}}_{al}^{\psi}(U)\oplus{\mathfrak{D}}_{al}^{\psi}(V)$. Using Prop. 3.7 $iii)$ and the fact that $\gamma_{1}^{U,V}({\mathsf{F}})+\gamma_{2}^{U,V}({\mathsf{F}})={\mathsf{F}}$ and $\gamma_{1}^{U^{\prime},V^{\prime}}({\mathsf{H}})+\gamma_{2}^{U^{\prime},V^{\prime}}({\mathsf{H}})={\mathsf{H}}$, we get $\tilde{\gamma}_{1}({\mathsf{F}})+\tilde{\gamma}_{2}({\mathsf{F}})={\mathsf{F}}$, as desired. ∎
We showed that one can attach Fréchet-Lie algebras ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}^{\psi}(U)$ to any $U\subset{\mathbb{R}}^{n}$. These form a pre-cosheaf over the pre-ordered set of subsets of ${\mathbb{R}}^{n}$. Moreover, Proposition 3.5, Corollary 3.1, and (when $\psi$ is gapped) Proposition 3.8 show that these spaces satisfy the cosheaf condition and Property I for transverse pairs $U,V\subset{\mathbb{R}}^{n}$. What prevents the functors ${\mathfrak{D}}_{al}$ and ${\mathfrak{D}}^{\psi}_{al}$ from forming local Lie algebras is the fact that pairs of subsets $U,V\subset{\mathbb{R}}^{n}$ generally do not intersect transversely. This problem can be resolved by restricting to a suitable set of subsets of ${\mathbb{R}}^{n}$ that have well-behaved intersections. In the next section we identify one such set, the set of semilinear subsets, prove that they form a Grothendieck site, and discuss some properties of this site.
4 The site of fuzzy semilinear sets
4.1 Semilinear sets and their thickenings
A semilinear set in ${\mathbb{R}}^{n}$ is a subset of ${\mathbb{R}}^{n}$ which can be defined by means of a finite number of linear equalities and strict linear inequalities. More precisely, a basic semilinear set in ${\mathbb{R}}^{n}$ is an intersection of a finite number of hyperplanes and open half-spaces, and a semilinear set is a finite union of basic semilinear sets. The set of semilinear subsets of ${\mathbb{R}}^{n}$ will be denoted ${\mathcal{S}}_{n}$. Projections ${\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ map ${\mathcal{S}}_{m+n}$ to ${\mathcal{S}}_{m}$ [23]. A map ${\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ is called semilinear iff its graph is a semilinear subset of ${\mathbb{R}}^{m+n}$. The composition of two semilinear maps is a semilinear map [23].
Recall that we use the $\ell^{\infty}$ metric on ${\mathbb{R}}^{n}$.
Lemma 4.1.
The distance function $d:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is semilinear.
Proof.
The function $(x,y)\mapsto x_{i}-y_{i}$ is semilinear for any $i$. The function $|\cdot|:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is semilinear. If $f,g:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ are semilinear, then $h=\max(f,g):{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is semilinear. Since the set of semilinear functions is closed under composition, this proves the lemma. ∎
Recall that for any set $U\subset{\mathbb{R}}^{n}$, we write $U^{r}:=\{x\in{\mathbb{R}}^{n}:\exists y\in U\text{ s.t }d(x,y)\leq r\}$ and call this the $r$-thickening of $U$. It is easy to see that if $U$ is closed, then $U^{r}$ is also closed for any $r\geq 0$.
Lemma 4.2.
If $U$ is semilinear, $U^{r}$ is semilinear for any $r$.
Proof.
Consider the set
$\Delta_{r}=\{(x,y)\in{\mathbb{R}}^{2n}|d(x,y)\leq r\}.$
By the previous lemma, $\Delta_{r}$ is semilinear. On the other hand, $U^{r}$ is the projection to the first ${\mathbb{R}}^{n}$ of $\Delta_{r}\cap({\mathbb{R}}^{n}\times U)\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}$. Since intersection and projection preserve the set of semilinear sets, the lemma is proved. ∎
Lemma 4.3.
If $U$ is convex, then $U^{r}$ is convex, for any $r\geq 0$.
Proof.
Suppose $x,y\in U^{r}$, and suppose $x^{\prime},y^{\prime}\in U$ satisfy $d(x,x^{\prime})\leq r$ and $d(y,y^{\prime})\leq r$. Then for any $t\in[0,1]$ we have $d(tx+(1-t)y,tx^{\prime}+(1-t)x^{\prime})\leq t^{2}d(x,x^{\prime})+(1-t)^{2}d(y,y^{\prime})\leq r.$
A polyhedron in ${\mathbb{R}}^{n}$ is an intersection of a finite number of closed half-spaces. A polyhedron is closed, but not necessarily compact. A closed semilinear set is the same as a finite union of polyhedra. Conversely, according to Theorem 19.6 from [24], a polyhedron can be described as a closed convex semilinear set. Combining this with the above lemmas, we get
Corollary 4.1.
If $U$ is a polyhedron, then $U^{r}$ is a polyhedron, for any $r\geq 0$.
4.2 A category of fuzzy semilinear sets
Clearly, if for $X,Y\in{\mathcal{S}}_{n}$ we have $X\subseteq Y$, then for any $r\geq 0$ we have $X^{r}\subseteq Y^{r}$. Also, for any $r,s\geq 0$ and any $U\in{\mathcal{S}}_{n}$ we have $(U^{r})^{s}\subseteq U^{r+s}$. Thus we can define a pre-order $\leq$ on ${\mathcal{S}}_{n}$ by saying that $U\leq V$ iff there exists $r\geq 0$ such that $U\subseteq V^{r}$. We will call this pre-order relation fuzzy inclusion. Equivalently, ${\mathcal{S}}_{n}$ can be made into a category, with a single morphism from $U$ to $V$ iff $U\leq V$. One can turn the pre-ordered set $({\mathcal{S}}_{n},\leq)$ into a poset by identifying isomorphic objects of the corresponding category, but for our purposes it is more convenient not to do so. On the other hand, every semilinear set is isomorphic to its closure, and we find it convenient to work with an equivalent category (or pre-ordered set) which contains only closed semilinear subsets. We will denote it ${\mathcal{C}\mathcal{S}}_{n}$ and call it the category (or pre-ordered set) of fuzzy semilinear sets.
Proposition 4.1.
${\mathcal{C}\mathcal{S}}_{n}$ has all pairwise joins: for any $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ the join is given by $U\cup V$.
Proof.
If $U\subseteq W^{r}$ and $V\subseteq W^{s}$ for some $r,s\geq 0$, then $U\subseteq W^{\max(r,s)}$ and $V\subseteq W^{\max(r,s)}$, and thus $U\cup V\subseteq W^{\max(r,s)}$. ∎
Let us show that ${\mathcal{C}\mathcal{S}}_{n}$ has all pairwise meets, using the notion of transverse intersection from the previous section. Recall (Definition 3.5) that we say two sets $U,V\subset{\mathbb{R}}^{n}$ are transverse if for some $C>0$ we have $d(x,U\cap V)\leq C\max(d(x,U),d(x,V))$ for all $x\in{\mathbb{R}}^{n}$.
We need the following geometric result [25] [^9] The proof in [25] is for the Euclidean distance, but since the Euclidean distance function and $d(x,y)=\|x-y\|_{\infty}$ are equivalent (each one is upper-bounded by a multiple of the other), the result applies to $d(x,y)$ as well. :
Lemma 4.4.
Let $P$ and $Q$ be polyhedra ${\mathbb{R}}^{n}$. If $P\cap Q\neq\varnothing$ then $P$ and $Q$ are transverse.
Proposition 4.2.
For every $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ there is an $r>0$ such that $U^{r}$ and $V^{r}$ are transverse.
Proof.
Let $U=\cup_{i}P_{i}$ and $V=\cup_{j}P_{j}$ be a decomposition of $U$ and $V$ into finite unions of polyhedra and choose $r>0$ so that $P_{i}^{r}\cup Q_{j}^{r}$ is nonempty for each pair $i,j$. By Corollary 4.1 and Lemma 4.4 there are constants $C_{P_{i}Q_{j}}>0$ such that
$d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$
for each pair $i,j$. Since $U^{r}=\cup_{i}P_{i}^{r}$ and $V^{r}=\cup_{i}Q_{i}^{r}$, for any $x\in{\mathbb{R}}$ there are indices $i^{*}$ and $j^{*}$ such that $d(x,U^{r})=d(x,P_{i^{*}}^{r})$ and $d(x,V^{r})=d(x,Q_{j^{*}}^{r})$. Then we have
$d(x,U^{r}\cap V^{r})$ $=d(x,\cup_{ij}(P_{i}^{r}\cap Q_{j}^{r}))$
$\leq d(x,P_{i^{*}}^{r}\cap Q_{j^{*}}^{r}))$
$\leq C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,P_{i^{*}}),d(x,Q_{j^{*}}))$
$=C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,U),d(x,V)),$
and so $U^{r},V^{r}$ are $C$-transverse for $C:=\max_{i,j}C_{P_{i}Q_{j}}$. ∎
Corollary 4.2.
With $U,V,r$ as above, $U^{r}\cap V^{r}$ is a meet of $U$ and $V$. In particular, ${\mathcal{C}\mathcal{S}}_{n}$ has all pairwise meets.
Proof.
Since $U$ and $U^{r}$ (resp. $V$ and $V^{r}$) are isomorphic in ${\mathcal{C}\mathcal{S}}_{n}$, it suffices to show that $U^{r}\cap V^{r}$ is a meet of $U^{r}$ and $V^{r}$. It’s clear that $U^{r}\cap V^{r}\leq U^{r}$ and $U^{r}\cap V^{r}\leq V^{r}$. Now suppose $W\in{\mathcal{C}\mathcal{S}}_{n}$ satisfies $W\leq U^{r}$ and $W\leq V^{r}$. Then there is an $s>0$ such that every $x\in W$ satisfies $\max(d(x,U^{r}),d(x,V^{r}))\leq s$. Since $d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$ for some $C>0$ we have $W\subset(U^{r}\cap V^{r})^{Cs}$. ∎
Proposition 4.3.
The pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$ is distributive.
Proof.
We need to show that for any $U,V,W\in{\mathcal{C}\mathcal{S}}_{n}$ we have $U\wedge(V\vee W)\leq(U\wedge V)\vee(U\wedge W)$. According to Corollary 4.2, there exists $r\geq 0$ such that $U\wedge(V\vee W)\simeq U^{r}\cap(V\cup W)^{r}$. Since $(V\cup W)^{r}=V^{r}\cup W^{r},$ we also have $U\wedge(V\vee W)\simeq(U^{r}\cap V^{r})\cup(U^{r}\cap W^{r})$. On the other hand, $U\wedge V\simeq U^{r}\wedge V^{r}\simeq(U^{r})^{s}\cap(V^{r})^{s}$ for some $s\geq 0$, and $U\wedge W\simeq U^{r}\wedge W^{r}\simeq(U^{r})^{t}\cap(W^{r})^{t}$ for some $t\geq 0$. Thus $(U\wedge V)\vee(U\wedge W)\simeq\left((U^{r})^{s}\cap(V^{r})^{s}\right)\cup\left((U^{r})^{t}\cap(W^{r})^{t}\right)$ for some $s,t\geq 0$. Since we have inclusions $U^{r}\cap V^{r}\subseteq(U^{r})^{s}\cap(V^{r})^{s}$ and $U^{r}\cap W^{r}\subseteq(U^{r})^{t}\cap(W^{r})^{t}$, the lemma is proved. ∎
We can now equip ${\mathcal{C}\mathcal{S}}_{n}$ with a Grothendieck topology of Section 2. There are two versions of it which differ in whether we allow empty covers of an initial object or not. In the case of the pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$, every bounded closed semilinear set is an initial object (they are all isomorphic objects of the category ${\mathcal{C}\mathcal{S}}_{n}$). Since such sets are not empty, we will disallow empty covers. This choice is also forced on us if we want certain pre-cosheaves to be cosheaves (see below). Note that we only consider non-empty closed semilinear sets.
More generally, for any $W\in{\mathcal{C}\mathcal{S}}_{n}$ we may consider a full sub-category ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ whose objects are $U\in{\mathcal{C}\mathcal{S}}_{n}$ such that $U\leq W$. This is a distributive pre-ordered set, and we will also have occasion to consider local Lie algebras on the associated site.
4.3 Spherical CS sets
Every two bounded elements of ${\mathcal{C}\mathcal{S}}_{n}$ are isomorphic objects of the corresponding category. More generally, any two elements of ${\mathcal{C}\mathcal{S}}_{n}$ which coincide outside some ball in ${\mathbb{R}}^{n}$ are isomorphic objects. Thus ${\mathcal{C}\mathcal{S}}_{n}$ encodes the large-scale structure of ${\mathbb{R}}^{n}$. To make this explicit, we will show that the pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$ is equivalent as a category to a certain poset of subsets of the “sphere at infinity” $S^{n-1}$.
A cone in ${\mathbb{R}}^{n}$ is a non-empty subset of ${\mathbb{R}}^{n}$ which is invariant under $x\mapsto\lambda x$, where $\lambda\geq 0$. Every cone contains the origin 00. Cones in ${\mathbb{R}}^{n}$ are in bijection with subsets of $S^{n-1}=({\mathbb{R}}^{n}\backslash\{0\})/{\mathbb{R}}^{*}_{+}$ where ${\mathbb{R}}^{*}_{+}$ is the group of positive real numbers under multiplication. If $A\subset S^{n-1}$, we denote the corresponding cone $c(A)$. In particular, $c(\varnothing)=\{0\}\in{\mathbb{R}}^{n}$. If $K$ is a cone in ${\mathbb{R}}^{n}$, we will denote the corresponding subset of $S^{n-1}$ by $\hat{K}$.
Definition 4.1.
$A\subset S^{n-1}$ is a spherical polyhedron iff $c(A)$ is a polyhedron and $A$ is contained in some open hemisphere of $S^{n-1}$. $A\subseteq S^{n-1}$ is a spherical CS set iff it is a union of a finite number of spherical polyhedra. The set of spherical CS sets in $S^{n-1}$ is denoted ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$.
Every polyhedron is convex and thus contractible. This implies:
Proposition 4.4.
Any spherical polyhedron is contractible.
Proof.
Without loss of generality, we may assume that the spherical polyhedron is contained in the hemisphere $S^{n-1}_{+}=S^{n-1}\cap\{x_{n}>0\}$. The map ${\mathbb{R}}^{n-1}\rightarrow S^{n-1}_{+}$ which sends $(x_{1},\ldots,x_{n-1})$ to the equivalence class of $(x_{1},\ldots,x_{n-1},1)$ is a homeomorphism which establishes a bijection between bounded polyhedra in ${\mathbb{R}}^{n-1}$ and spherical polyhedra contained in $S^{n-1}_{+}$. ∎
Proposition 4.5.
The intersection of two spherical polyhedra is a spherical polyhedron. The union and intersection of two spherical CS sets is a spherical CS set.
Proof.
Clear from definitions. ∎
Spherical CS sets form a poset ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ under inclusion. This poset has pairwise joins and meets given by unions and intersections, respectively.
Proposition 4.6.
The category ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ is equivalent to the category ${\mathcal{C}\mathcal{S}}_{n}$.
Proof.
We proceed first by defining a functor from ${\mathcal{C}\mathcal{S}}_{n}$ to ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. Notice, it is sufficient to define a functor on isomorphism classes of polyhedrons then extend by joins. Every closed polyhedron is a finite intersection of closed half spaces $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\},$ where for each $i=1,\ldots k$, $n_{i}\in{\mathbb{R}}^{n}$ is the unit normal vector of the $i$th supporting hyperplane and $b_{i}\in{\mathbb{R}}$. By definition, $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\}$ is isomorphic to $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq 0\}$ in ${\mathcal{C}\mathcal{S}}_{n}$. In other words, every polyhedron is isomorphic to a cone $\{Nx\preceq 0\},$ where $N$ is the matrix with $n_{i}$ as rows and we write $y\preceq 0$ for $y\in{\mathbb{R}}^{n}$ when $y_{i}\leq 0$ for all $i=1,\ldots n$. If $\{n_{i}\}_{i=1}^{k}$ spans ${\mathbb{R}}^{n}$, the set $\{Nx\preceq 0\}$, which contains no antipodal points, is mapped to its corresponding spherical polyhedron. Otherwise, complete $\{n_{i}\}_{i=1}^{k}$ into a spanning set $\{n_{1},\dots,n_{k},m_{1},\dots,m_{l}\}$ and decompose $\{Nx\preceq 0\}$ into a union of cones $\{Nx\preceq 0\}\cap\bigcap_{j}\{\pm m_{j}\cdot x\leq 0\}$. By the universal property of the join, an arbitrary closed semilinear set is isomorphic to a union of conical polyhedra. For functoriality, it is sufficient to note that for any pair of closed cones $U,V$ a fuzzy inclusion $U\subseteq V^{r}$ for some $r\geq 0$ implies $U\subseteq V$. It is easy to check that this functor is fully faithful and (essentially) surjective. ∎
Note that under this equivalence all bounded elements of ${\mathcal{C}\mathcal{S}}_{n}$ correspond to $\varnothing\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. The canonical Grothendieck topology on ${\mathcal{C}\mathcal{S}}_{n}$ corresponds to a slightly unusual Grothendieck topology on the poset of spherical CS subsets of $S^{n-1}$: the one where empty covers of $\varnothing$ are not allowed. Consequently, a cosheaf of vector spaces on ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ equipped with this topology need not map $\varnothing$ to the zero vector space.
4.4 Spherical CS cohomology
Let ${\mathfrak{U}}$ be a spherical CS cover of a spherical CS set $A$. Spherical CS cohomology $\check{H}^{\bullet}_{CS}({\mathfrak{U}},A;{\mathbb{R}})$ is defined to be the simplicial cohomology of the Čech nerve $N({\mathfrak{U}})$.
Definition 4.2.
[…]
The following proposition connects spherical CS cohomology with singular cohomology using a functorial version of nerve theorems [26, 27].
Proposition 4.7.
For any spherical CS set $A$ the graded vector space $\check{H}^{\bullet}_{CS}(A,{\mathbb{R}})$ is isomorphic to the singular cohomology $H^{\bullet}(A,{\mathbb{R}})$.
Proof.
[…]
5 Local Lie algebras over fuzzy semilinear sets
5.1 Basic examples
Let $\psi$ be a state of a quantum lattice system on ${\mathbb{R}}^{n}$. By Lemma 3.2 the maps sending $U\in{\mathcal{C}\mathcal{S}}_{n}$ to ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}^{\psi}(U)$ are pre-cosheaves of Fréchet spaces on ${\mathcal{C}\mathcal{S}}_{n}$. We denote these pre-cosheaves by ${\mathfrak{D}}_{al}$ and ${\mathfrak{D}}_{al}^{\psi}$. Putting together Lemma 2.1, Proposition 3.5, Corollary 3.1, and Proposition 3.8, we have
Theorem 5.1.
${\mathfrak{D}}_{al}$ is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. If $\psi$ is gapped, then ${\mathfrak{D}}_{al}^{\psi}$ is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$.
Our main object of study is the local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ attached to a gapped state $\psi$ of a lattice system $(\Lambda,\{V_{j}\}_{j\in\Lambda})$.
A much simpler example of a local Lie algebra arises from a finite-dimensional Lie algebra ${\mathfrak{g}}$ and any subset $\Lambda\subset{\mathbb{R}}^{n}$. For any $U\in{\mathcal{C}\mathcal{S}}_{n}$ let ${{\mathfrak{g}}_{al}}(U)$ be the space of bounded functions $U\cap\Lambda\rightarrow{\mathfrak{g}}$ which decay superpolynomially away from $U$. The subscript “al” stands for “almost localized”. More precisely, ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})$ is the space of bounded functions $\Lambda\rightarrow{\mathfrak{g}}$, while ${{\mathfrak{g}}_{al}}(U)$ is defined as a subspace of ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})$ consisting of functions $f:\Lambda\rightarrow{\mathfrak{g}}$ such that the following seminorms are finite:
$p_{k,U}(f)=\sup_{j\in\Lambda}|f(j)|(1+d(U,j))^{k},\quad k\in{\mathbb{N}}.$
Proposition 5.1.
The assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ is a local Lie algebra.
Proof.
It is easy to check that the assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ it is a coflasque pre-cosheaf of Fréchet-Lie algebras satisfying Property I. The only thing left to check is that it is a cosheaf of vector spaces. By Lemma 2.1, it is sufficient to show that for any $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ the sequence
${{\mathfrak{g}}_{al}}(U\wedge V)\rightarrow{{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)\rightarrow{{\mathfrak{g}}_{al}}(U\vee V)\rightarrow 0$
is exact. To show exactness at ${{\mathfrak{g}}_{al}}(U\vee V)={{\mathfrak{g}}_{al}}(U\cup V)$, we note that every $f\in{{\mathfrak{g}}_{al}}(U\cup V)$ can be written as a sum $f_{U}+f_{V}$, where
$f_{U}(x)=\begin{cases}f(x),&d(x,U)<d(x,V),\\ \frac{1}{2}f(x),&d(x,U)=d(x,V),\\ 0,&d(x,U)>d(x,V),\end{cases}$
and $f_{V}(x)$ is defined by a similar expression with $U$ and $V$ exchanged. Using $d(x,U\cup V)=\min(d(x,U),d(x,V))$ it is easy to check that $p_{k,U}(f_{U})\leq p_{k,U\cup V}(f)$ and $p_{k,V}(f_{V})\leq p_{k,U\cup V}(f)$ for all $k$, and thus $f_{U}\in{{\mathfrak{g}}_{al}}(U)$ and $f_{V}\in{{\mathfrak{g}}_{al}}(V)$.
To show exactness at ${{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)$, suppose $f\in{{\mathfrak{g}}_{al}}(U)\cap{{\mathfrak{g}}_{al}}(V)$. From the proof of Prop. 4.2, there exist $r\geq 0,C_{UV}>0$ such that $d(x,U^{r}\cap V^{r})\leq C_{UV}\max(d(x,U),d(x,V))$. We may assume that $C_{UV}\geq 1$, in which case for any $x\in{\mathbb{R}}^{n}$ and any $k\in{\mathbb{N}}$
$(1+d(x,U^{r}\cap V^{r}))^{k}\leq C^{k}_{UV}(1+\max(d(x,U),d(x,V))^{k}.$
Therefore for any $f\in{{\mathfrak{g}}_{al}}(U)\cap{{\mathfrak{g}}_{al}}(V)$ we have
$p_{k,U^{r}\cap V^{r}}(f)\leq C^{k}_{UV}\max(p_{k,U}(f),p_{k,V}(f)).$
Since by Corollary 4.2 one can take $U\wedge V=U^{r}\cap V^{r}$, we conclude that $f\in{{\mathfrak{g}}_{al}}(U\wedge V)$. ∎
5.2 Symmetries of lattice systems
If $\Lambda\subset{\mathbb{R}}^{n}$ is countable, it can be viewed as a lattice in the physical sense, and the local Lie algebra ${{\mathfrak{g}}_{al}}$ over ${\mathcal{C}\mathcal{S}}_{n}$ models infinitesimal gauge transformations of a lattice system on ${\mathbb{R}}^{n}$.
Definition 5.1.
A local action of a compact Lie group on a lattice system $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ is a collection of homomorphisms $\rho_{j}:G\rightarrow U(V_{j})$ such that the norms of the corresponding Lie algebra homomorphisms ${\mathfrak{g}}\rightarrow B(V_{j})$ are bounded uniformly in $j$.
A local action of $G$ on $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ gives rise to a homomorphism from $G$ to the automorphism group of ${\mathscr{A}}$ via $g\mapsto\otimes_{j\in\Lambda}{\rm Ad}_{\rho_{j}(g)}$ which is smooth on ${\mathscr{A}}_{a\ell}$. The corresponding generator ${\mathsf{Q}}$ is a homomorphism from ${\mathfrak{g}}$ to the Lie algebra of derivations of ${\mathscr{A}}_{a\ell}$ defined by
${\mathsf{Q}}:(a,{\mathcal{A}})\mapsto\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(a),{\mathcal{A}}],\quad a\in{\mathfrak{g}},\ {\mathcal{A}}\in{\mathscr{A}}_{a\ell},$
where ${\mathsf{q}}_{j}$ is the traceless part of the generator of $\rho_{j}$. The image of ${\mathsf{Q}}$ lands in ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$, with ${\mathsf{Q}}(a)^{Y}=\sum_{j\in\Lambda}{\mathsf{q}}_{j}(a)^{Y}$. We will regard ${\mathsf{Q}}$ as a homomorphism of Fréchet-Lie algebras ${\mathfrak{g}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$.
This morphism of Fréchet-Lie algebras can be lifted to a morphism of local Lie algebras ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. Indeed, for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ and any $f\in{{\mathfrak{g}}_{al}}(U)$ we let ${\mathsf{Q}}(f)$ be a derivation of ${\mathscr{A}}_{a\ell}$ given by
${\mathsf{Q}}(f)({\mathcal{A}})=\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(f(j)),{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}}_{a\ell}.$
It is easy to check that this derivation belongs to ${\mathfrak{D}}_{al}(U)$ and that the above map is a continuous homomorphism ${{\mathfrak{g}}_{al}}(U)\rightarrow{\mathfrak{D}}_{al}(U)$. The physical interpretation is that a local action of a compact Lie group on a quantum lattice system can be gauged on the infinitesimal level.
Definition 5.2.
A state $\psi$ of ${\mathscr{A}}$ is said to be invariant under a local action of a compact Lie group $G$ if it is invariant under the corresponding automorphisms of ${\mathscr{A}}$.
Let $\psi$ be a gapped state of ${\mathscr{A}}$ invariant under a local action of a Lie group $G$. In that case the image of ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ lands in ${\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. One may ask if this morphism of Fréchet-Lie algebras can be lifted to a morphism of local Lie algebras ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$. If this is the case, then the symmetry $G$ of $\psi$ can be gauged on the infinitesimal level. In the next section we construct obstructions for the existence of such a morphism of local Lie algebras and show that zero-temperature Hall conductance is an example of such an obstruction.
5.3 Equivariantization
As a preliminary step, for any $G$-invariant gapped state $\psi$ we are going to define a graded local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ which is a $G$-equivariant version of the local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$. Recall that a graded local Lie algebra is a cosheaf of graded vector spaces that is a pre-cosheaf of graded Lie algebras satisfying the graded analogue of Property I. For example, if ${\mathfrak{F}}$ is a local Lie algebra and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally finite supercommutative graded algebra with finite-dimensional graded factors $A_{k}$, [^10] A graded vector space is locally finite iff its graded components are finite-dimensional. then $U\mapsto{\mathfrak{F}}(U)\otimes A$ is a graded local Lie algebra. We denote it ${\mathfrak{F}}\otimes A$.
Fix a compact Lie group $G$ and a distributive pre-ordered set $X$ and consider the category of graded local Lie algebras over $X$ equipped with a $G$-action. An object of this category is a graded local Lie algebra ${\mathfrak{F}}$ on which $G$ acts by automorphisms; morphisms are defined in an obvious manner. The first step is to define a functor ${\mathfrak{F}}\mapsto{\mathfrak{F}}^{G}$ from this category to the category of graded local Lie algebras over $X$ such that ${\mathfrak{F}}^{G}(U)$ is the Lie algebra of $G$-invariant elements of ${\mathfrak{F}}(U)$. It is clear how to define such a functor for coflasque pre-cosheaves of Lie algebras with Property I, but the pre-cosheaf ${\mathfrak{F}}^{G}$ will not be a cosheaf of vector spaces without further assumptions about ${\mathfrak{F}}$ and the $G$-action.
Definition 5.3.
An action of $G$ on a pre-cosheaf of Fréchet spaces ${\mathfrak{F}}$ is smooth if for each $U\in X$ the seminorms defining the topology of ${\mathfrak{F}}(U)$ can be chosen to be $G$-invariant and the map $G\times{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(U)$ defining the action is smooth. An action of $G$ on a pre-cosheaf of graded Fréchet spaces is smooth is the $G$-action on every graded component is smooth.
Proposition 5.2.
Let ${\mathfrak{F}}$ be a pre-cosheaf of graded Fréchet  spaces over $X$ equipped with a smooth action of a compact Lie group $G$. The assignment $U\mapsto{\mathfrak{F}}^{G}(U)=\left({\mathfrak{F}}(U)\right)^{G}$ is a cosheaf of graded Fréchet spaces.
Proof.
It is sufficient to prove this in the ungraded case. We need to show that for any $U,V\in X$ the sequence
[…]
is exact. To show exactness at the rightmost term, let ${\mathsf{F}}$ be a $G$-invariant element of ${\mathfrak{F}}(U\vee V)$ and let ${\mathsf{F}}_{U}\in{\mathfrak{F}}(U)$ and ${\mathsf{F}}_{V}\in{\mathfrak{F}}(V)$ be such that $\iota_{U\cup V,U}{\mathsf{F}}_{U}+\iota_{U\cup V,V}{\mathsf{F}}_{V}={\mathsf{F}}$. Averaging the action map $G\times{\mathfrak{F}}(U)$ over $G$ with the Haar measure gives a linear map $h_{U}:{\mathfrak{F}}(U)\mapsto{\mathfrak{F}}^{G}(U)$ which is identity when restricted to ${\mathfrak{F}}^{G}(U)$. The co-restriction morphisms intertwine these maps. Thus $\iota_{U\cup V,U}\circ h_{U}({\mathsf{F}}_{U})+\iota_{U\cup V,V}\circ h_{V}({\mathsf{F}}_{V})={\mathsf{F}}$ which proves that $\beta$ is surjective. Exactness in the middle term is proved similarly. ∎
Remark 5.1.
For the proof to go through, it suffices to require the map $G\times{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(U)$ to be continuous. However, if it is smooth, ${\mathfrak{F}}(U)$ becomes a ${\mathfrak{g}}$-module and all elements in ${\mathfrak{F}}^{G}(U)\subset{\mathfrak{F}}(U)$ are annihilated by the ${\mathfrak{g}}$-action. We will use this later on.
Example 5.1.
If $G$ acts locally on a lattice system, the action of $G$ on the local Lie algebra ${\mathfrak{D}}_{al}$ is smooth. If $\psi$ is a $G$-invariant gapped state of such a lattice system, the action of $G$ on ${\mathfrak{D}}^{\psi}_{al}$ is smooth.
Example 5.2.
Consider the local Lie algebra ${{\mathfrak{g}}_{al}}$ over ${\mathcal{C}\mathcal{S}}_{n}$ associated to a finite-dimensional Lie algebra ${\mathfrak{g}}$ (see Section 4). Assume that ${\mathfrak{g}}$ is the Lie algebra of a compact Lie group $G$, then there is an obvious $G$-action on ${{\mathfrak{g}}_{al}}$:
$(g\cdot f)(j)={\rm Ad}_{g}f(j),\quad g\in G,j\in\Lambda.$
This $G$-action is smooth.
Example 5.3.
If ${\mathfrak{F}}$ is a local Fréchet-Lie algebra with a smooth $G$-action and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally-finite supercommutative graded algebra on which $G$ acts by automorphisms, then the $G$-action on ${\mathfrak{F}}\otimes A$ is smooth.
Corollary 5.1.
Let ${\mathfrak{F}}$ be a graded local Fréchet-Lie algebra over $X$ with a smooth $G$-action. The functor of $G$-invariant elements maps ${\mathfrak{F}}$ to a graded local FréchetLie algebra ${\mathfrak{F}}^{G}$ over $X$.
Definition 5.4.
Let ${\mathfrak{F}}$ be a local Fréchet-Lie algebra over $X$ with a smooth $G$-action. The $G$-equivariantization functor sends ${\mathfrak{F}}$ to the negatively-graded local Fréchet-Lie algebra ${\mathfrak{F}}^{\bf G}$ defined by
$U\mapsto\left({\mathfrak{F}}(U)\otimes\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])\right)^{G}.$
In the cases of interest to us, the $G$-action on a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ is infinitesimally inner, in the sense that the ${\mathfrak{g}}$-module structure mentioned in Remark 5.1 arises from a homomorphism $\rho:{\mathfrak{g}}\rightarrow{\mathfrak{F}}({\mathbb{R}}^{n})$. In such a case, the graded local Lie algebra ${\mathfrak{F}}^{\bf G}$ has an extra bit of structure: a central element in ${\mathfrak{F}}^{\bf G}({\mathbb{R}}^{n})$ of degree $-2$. This element is simply $\rho$ re-interpreted as an element of ${\mathfrak{F}}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$. In the terminology of Section 2, ${\mathfrak{F}}^{\bf G}$ is a pointed graded local Fréchet-Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. It is easy to see that the $G$-equivariantization functor respects this extra structure. That is, if $f:{\mathfrak{F}}\rightarrow{\mathfrak{F}}^{\prime}$ is a morphism of local Fréchet-Lie algebras over ${\mathcal{C}\mathcal{S}}_{n}$ commuting with infinitesimally inner smooth $G$-actions on ${\mathfrak{F}}$ and ${\mathfrak{F}}^{\prime}$, then $f^{\bf G}$ maps the central element $\rho\in{\mathfrak{F}}^{\bf G}({\mathbb{R}}^{n})_{-2}$ to the central element $\rho^{\prime}\in{{\mathfrak{F}}^{\prime}}^{\bf G}({\mathbb{R}}^{n})_{-2}$.
Example 5.4.
Let $\psi$ be a $G$-invariant gapped state of a quantum lattice system with a local $G$-action which on the infinitesimal level is described by ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Consider the graded local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ with its smooth $G$-action (Example 5.1) and its $G$-equivariantization ${\mathfrak{D}}^{\psi,\bf G}_{al}$. The distinguished central element of ${\mathfrak{D}}^{\psi,\bf G}_{al}({\mathbb{R}}^{n})$ is ${\mathsf{Q}}$ regarded as an element of ${\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$.
Example 5.5.
Consider the graded local Lie algebra ${{\mathfrak{g}}^{\bf G}_{al}}$ of Example 5.2. The degree $-2$ component of ${{\mathfrak{g}}^{\bf G}_{al}}({\mathbb{R}}^{n})$ is the space of $G$-invariant bounded functions on $\Lambda$ with values in ${\mathfrak{g}}\otimes{\mathfrak{g}}^{*}$. The distinguished central element is the constant function on $\Lambda$ which takes the value ${\rm id}_{\mathfrak{g}}$.
Armed with the equivariantization functor, we can now explain our strategy for constructing obstructions for the existence of a local Lie algebra morphism ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ which lifts the Fréchet-Lie algebra morphism ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Suppose such a morphism $\rho$ exists. Applying the $G$-equivariantization functor, we get a morphism of pointed negatively-graded local Fréchet-Lie algebras $\rho^{\bf G}:{{\mathfrak{g}}^{\bf G}_{al}}\rightarrow{\mathfrak{D}}^{\psi,\bf G}_{al}$. For any CS cover ${\mathfrak{U}}$ of ${\mathbb{R}}^{n}$ an application of the Čech functor gives a morphism of acyclic pointed DGLAs
$C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{{\mathfrak{g}}^{\bf G}_{al}})\rightarrow C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al}).$
Consequently, a obstruction for the existence of such a pointed DGLA morphism is an obstruction for the existence of $\rho$. In the next section we use the twisted Maurer-Cartan equation for pointed DGLAs to construct such obstructions and identify them as topological invariants of gapped states defined in [11].
6 Topological invariants of $G$-invariant gapped states
6.1 The commutator class
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pointed DGLA, i.e. a DGLA with a distinguished central cycle ${\mathsf{B}}\in{\mathfrak{M}}_{-2}$. Assume it is a limit of an inverse system of nilpotent pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$.
Definition 6.1.
The commutator DGLA, denoted by ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$, is defined to be the closure of the commutator subalgebra of ${\mathfrak{M}}$. Namely, ${\mathsf{q}}\in{\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ if and only if for any $N$ its projection to ${\mathfrak{M}}_{N}$ is a finite linear combination of commutators in ${\mathfrak{M}}_{N}$.
Even if ${\mathfrak{M}}$ is acyclic, the DGLA ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ is not necessarily acyclic. We would like to construct an obstruction to finding an element ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ which satisfies $\partial{\mathsf{p}}={\mathsf{B}}$ and $[{\mathsf{p}},{\mathsf{p}}]=0$.
Definition 6.2.
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pointed DGLA. A ${\mathsf{B}}$-twisted Maurer-Cartan element in ${\mathfrak{M}}$ is ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ which satisfies
$\partial{\mathsf{p}}+\frac{1}{2}[{\mathsf{p}},{\mathsf{p}}]={\mathsf{B}}.$
We will denote the set of ${\mathsf{B}}$-twisted MC elements of ${\mathfrak{M}}$ by ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$. The map $({\mathfrak{M}},{\mathsf{B}})\mapsto{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ can be upgraded to a functor from the category of pointed DGLAs to the category of sets in an obvious way.
Let $({\mathfrak{M}},{\mathsf{B}})$ be pronilpotent pointed DGLA and ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then $[{\mathsf{p}},{\mathsf{p}}]$ is a cycle of the DGLA ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$.
Proposition 6.1.
Let $({\mathfrak{M}},{\mathsf{B}})$ be an acyclic pointed DGLA which is a limit of an inverse system of nilpotent acyclic pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. Assume further that the structure morphisms $r_{N,N-1}:{\mathfrak{M}}_{N}\rightarrow{\mathfrak{M}}_{N-1}$ are surjective and ${\mathfrak{j}}_{N}=\ker r_{N,N-1}$ is central in ${\mathfrak{M}}_{N}$. Finally, assume $[{\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$. Then ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. Furthermore, the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ in ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ for ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is independent of the choice of ${\mathsf{p}}$.
A proof of this result can be found in Appendix B. It uses some results from deformation theory. We will call the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ the commutator class of the acyclic pointed DGLA $({\mathfrak{M}},{\mathsf{B}})$, for lack of a better name, and denote it ${\mathbf{com}}({\mathfrak{M}},{\mathsf{B}})$. The assignment $({\mathfrak{M}},{\mathsf{B}})\mapsto(H_{\bullet}({\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}),{\mathbf{com}}({\mathfrak{M}},{\mathsf{B}}))$ is a functor from the full sub-category of the category of acyclic pointed DGLAs satisfying the conditions of Proposition 6.1, to the category ${{\mathsf{pVect}}_{\mathbb{Z}}}$ of pointed graded vector spaces.
The conditions of Prop. 6.1 apply to any pointed DGLA which is the value of the Čech functor on a graded local Lie algebra ${\mathfrak{F}}^{\bf G}$ over some $(X,\leq)$, where ${\mathfrak{F}}$ is a local Lie algebra equipped with a smooth infinitesimally inner $G$-action. Indeed, along with the graded algebra $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$ used in the construction of ${\mathfrak{F}}^{\bf G}$ one can consider its quotient by the ideal $J_{N}=\prod_{k=N+1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$. Replacing $A$ with the nilpotent graded algebra $A/J_{N}$ throughout, for any cover ${\mathfrak{U}}$ of the terminal object $T$ one gets a sequence of nilpotent acyclic pointed DGLAs labeled by $N\in{\mathbb{N}}$. They assemble into an inverse system in an obvious manner, and its limit is the acyclic pointed DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}}^{\bf G}).$ It is easy to see that the remaining conditions of Prop. 6.1 are also satisfied. In particular, Prop. 6.1 applies to the pointed DGLAs associated the graded local Lie algebras ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and ${{\mathfrak{g}}^{\bf G}_{al}}$ and any CS cover of ${\mathbb{R}}^{n}$.
Proposition 6.2.
For any pointed DGLA $({\mathfrak{M}}({\mathfrak{U}}),{\mathsf{B}})$ obtained, as above, from the Čech functor with respect to a cover ${\mathfrak{U}}$, ${\rm MC}$ is functorial in ${\mathfrak{U}}$.
Proof.
Let ${\mathfrak{V}}=\{V_{j}\}_{j\in J}$ be a refinement of ${\mathfrak{U}}$. By definition, there exists a map $\phi:J\rightarrow I$ with $V_{j}\subset U_{\phi(j)}.$ According to Prop. 2.2, there is a map from ${\mathfrak{M}}({\mathfrak{V}})$ to ${\mathfrak{M}}({\mathfrak{U}})$. As ${\mathsf{B}}$ is unaffected by $\phi_{*}$, we deduce from ${\mathsf{p}}^{\mathfrak{V}}\in{\rm MC}({\mathfrak{M}}({\mathfrak{V}}),{\mathsf{B}})$ that $\phi_{*}{\mathsf{p}}^{\mathfrak{V}}\in{\rm MC}({\mathfrak{M}}({\mathfrak{U}}),{\mathsf{B}})$. ∎
Example 6.1.
Let $G$ be a compact Lie group, ${\mathfrak{g}}$ be its Lie algebra, and ${{\mathfrak{g}}_{al}}$ be the pointed local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ of Examples (5.2) and (5.5). The distinguished central cycle in $C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n},{{\mathfrak{g}}^{\bf G}_{al}})$ is the constant function on $\Lambda$ with value ${\rm id}_{\mathfrak{g}}$ Here ${\rm id}_{\mathfrak{g}}$ is regarded as a $G$-invariant element of ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$. For any cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ one can construct a twisted MC element ${\mathsf{q}}$ as follows. Pick $r>0$ large enough so that the interiors of $U_{i}^{r}$, $i\in I$, cover ${\mathbb{R}}^{n}$ in the usual sense and pick a partition of unity $\chi_{i}$, $i\in I$, subordinate to this open cover. For any $j\in\Lambda$ and any $i\in I$ let ${\mathsf{q}}_{i}(j)=\chi_{i}(j){\rm id}_{\mathfrak{g}}$. It is easy to verify that this is a twisted MC element satisfying $[{\mathsf{q}},{\mathsf{q}}]=0$. Thus the commutator class vanishes in this case.
Let $G$ be a compact Lie group, $\psi$ be a gapped $G$-invariant state of a lattice system on ${\mathbb{R}}^{n}$, and $U\in{\mathcal{C}\mathcal{S}}_{n}$. The commutator class of the acyclic pointed DGLA $(C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al}),{\mathsf{Q}})$ is an obstruction for the existence of a morphism of local Lie algebras $\rho:{{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ which is a lift of the Lie algebra morphism ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Indeed, as explained in Section 5.3, if $\rho$ exists, it induces a morphism of pointed DGLAs $C^{aug}_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{{\mathfrak{g}}_{al}})\rightarrow C^{aug}_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ which in turn induces a morphism in the category ${{\mathsf{pVect}}_{\mathbb{Z}}}$ which maps the commutator class of ${{\mathfrak{g}}_{al}}$ to the commutator class of ${\mathfrak{D}}^{\psi,\bf G}_{al}$. Since the former class vanishes (see Example 6.1), so must the latter.
6.2 Construction of topological invariants
The commutator class defined in the previous section is not a useful invariant of a gapped $G$-invariant state $\psi$ because it takes values in a set which itself depends on $\psi$. It is also not invariant under LGAs (defined in Section 3.2) and thus is not an invariant of a gapped phase (see Remark 3.2). In this section we define a pairing between the commutator classes of ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and spherical CS cohomology classes of the sphere at infinity $S^{n-1}$, which takes values in the algebra of $G$-invariant symmetric polynomials on ${\mathfrak{g}}$. This gives a useful invariant of a gapped phase which is also an obstruction to promoting the global symmetry $G$ of the state to a local symmetry. Additionally, we will show that the invariant does not depends on the choice of the cover ${\mathfrak{U}}$ and thus is essentially unique.
Keeping in view further generalizations, we work over a sub-site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ [^11] This is the so-called overcategory whose objects are equipped with a morphism to $W$. Due to coflasqueness, this amounts to a restriction to objects fuzzily included in $W$. which depends on an arbitray $W\in{\mathcal{C}\mathcal{S}}_{n}$. Let ${\hat{W}}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ be the image of $W$ under the equivalence between ${\mathcal{C}\mathcal{S}}_{n}$ and ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. We continue to denote by ${\mathfrak{D}}_{al}$ the local Fréchet-Lie algebra which maps $U\in{{{\mathcal{C}\mathcal{S}}_{n}}/W}$ to ${\mathfrak{D}}_{al}$. For any cover ${\mathfrak{U}}$ of $W$ we denote by $\hat{\mathfrak{U}}$ the corresponding cover of $\hat{W}$.
Definition 6.3.
For any ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ covering $W\in{\mathcal{C}\mathcal{S}}_{n}$, any ${\mathsf{f}}\in C_{p}({\mathfrak{U}},W;{\mathfrak{D}}_{al})$, and any $\beta\in\check{C}^{p}(\hat{\mathfrak{U}},\hat{W};{\mathbb{R}})$ define an evaluation
$\langle{\mathsf{f}},\beta\rangle=\sum_{s\in I_{p}}\beta_{s}{\mathsf{f}}_{s}\in{\mathfrak{D}}_{al}(W),$
where $I_{k}:=\{i_{0}<i_{1}<\dots<i_{k}\in I^{k+1}\}$. We adopt the convention that $\beta_{s}=0$ for $\bigwedge_{j\in s}U_{j}$ bounded. The definition implicitly uses co-restriction of the coflasque cosheaf.
Lemma 6.1.
For any ${\mathfrak{U}}$ covering $W\in{\mathcal{C}\mathcal{S}}_{n}$, any cycle ${\mathsf{f}}\in C_{p}({\mathfrak{U}},W;{\mathfrak{D}}_{al})$, and any cocycle $\beta\in\check{C}^{p}(\hat{\mathfrak{U}},\hat{W};{\mathbb{R}})$ the derivation $\langle{\mathsf{f}},\beta\rangle\in{\mathfrak{D}}_{al}(W)$ is inner.
Proof.
The chain complex
$\ldots\rightarrow\bigoplus_{i<j}{\mathfrak{D}}_{al}(U_{i}\wedge U_{j})\rightarrow\bigoplus_{i}{\mathfrak{D}}_{al}(U_{i})\rightarrow{\mathfrak{D}}_{al}(W)\rightarrow 0$ (23)
is acyclic. Since ${\mathsf{f}}$ is a cycle, ${\mathsf{f}}=\partial{\mathsf{g}}$ for some ${\mathsf{g}}\in\bigoplus_{s\in I_{p+1}}{\mathfrak{D}}_{al}(\bigwedge_{j\in s}U_{j}).$ Thus
$\langle{\mathsf{f}},\beta\rangle=$ $\langle\partial{\mathsf{g}},\beta\rangle$
$=$ $\langle{\mathsf{g}},\partial^{*}\beta\rangle$
$=$ $\sum_{s\in I_{p+1}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s}$
$=$ $\sum_{\begin{subarray}{c}s\in I_{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$
where $\partial^{*}$ is the adjoint of $\partial$ defined by
$(\partial^{*}\beta)_{i_{0}\dots i_{p}}=\sum_{k=0}^{p}(-1)^{k}\beta_{i_{0}\dots\widehat{i_{k}}\dots i_{p}}.$
The last expression is clearly inner because each ${\mathsf{g}}_{s}$ is almost localized near some bounded region. The last equality depends on vanishing of
$\sum_{\begin{subarray}{c}s\in I_{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$
because when $\bigwedge_{j\in s}U_{j}$ is unbounded $(\partial^{*}\beta)_{s}=(\check{\delta}\beta)_{s}=0$ where $\check{\delta}$ is the coboundary of $\check{C}^{p}(\hat{\mathfrak{U}},\hat{W};{\mathbb{R}})$. ∎
Remark 6.1.
The above lemma remains true if one replaces ${\mathfrak{D}}_{al}$ with ${\mathfrak{D}}_{al}\otimes A$ where $A$ is any locally-finite graded vector space. Then the pairing $\langle{\mathsf{f}},\beta\rangle$ takes values in the space of inner derivations ${{\mathfrak{d}}_{al}}$ (Definition 3.3) tensored with $A$.
Remark 6.2.
The relation between $\partial^{*}$ and $\check{\delta}$ is as follows. They are equal for $s\in I_{p+1}$ with $\bigwedge_{j\in s}U_{j}$ unbounded. For $s$ with $\bigwedge_{j\in s}U_{j}$ bounded, $\bigcap_{j\in s}\hat{U}_{j}=\varnothing$ and $(\partial^{*}\beta)_{s}$ is in general nonzero whereas $(\check{\delta}\beta)_{s}=0$. This discrepancy arises because the Grothendieck topology on the poset of spherical CS sets induced by the coherent topology on ${\mathcal{C}\mathcal{S}}_{n}$ does not allow the empty cover of the empty set. It is this discrepancy that makes the evaluation of DGLA cycles on spherical CS cocycles non-trivial.
If $U,V$ are closed semilinear sets with a bounded meet then any derivation in $[{\mathfrak{D}}_{al}(U),{\mathfrak{D}}_{al}(V)]$ is inner and thus has a well-defined average in any state on ${\mathscr{A}}$. More generally, if $A$ is a locally-finite supercommutative graded algebra and ${\mathfrak{F}}={\mathfrak{D}}_{al}\otimes A$ or some sub-algebra thereof, any element of $[{\mathfrak{F}}(U),{\mathfrak{F}}(V)]$ for $U\wedge V$ bounded has a well-defined average in any state of ${\mathscr{A}}$. The average takes values in $A$. We will need the following more refined result:
Lemma 6.2.
Let $U$ and $V$ be closed semilinear sets such that $U\wedge V$ is bounded. Let $A$ be a locally-finite supercommutative graded algebra and $\psi$ be a state. Then $\psi([{\mathfrak{D}}^{\psi}_{al}(U)\otimes A,{\mathfrak{D}}_{al}(V)\otimes A])=0.$
Proof.
It suffices to consider the case $A={\mathbb{R}}$. Let ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U)$ and ${\mathsf{G}}\in{\mathfrak{D}}_{al}(V)$. Propositions 3.2 and 3.6 give
$\psi([{\mathsf{F}},{\mathsf{G}}])=\sum_{X\cap Y\neq\varnothing}\psi([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])=\sum_{Y}\psi({\mathsf{F}}({\mathsf{G}}^{Y}))=0,$
where the last equality is because ${\mathsf{F}}$ preserves $\psi$. ∎
In what follows we will apply Lemma 6.1 to the graded local Lie algebra ${\mathfrak{D}}^{\psi}_{al}\otimes A$ (see Remark 6.1). The evaluation of cycles of $C^{aug}_{\bullet}({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A)$ on spherical CS cocycles has especially nice properties when DGLA cycles belong to the commutator DGLA
$\overline{\left[C^{aug}_{\bullet+1}\left({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A\right),C^{aug}_{\bullet+1}\left({\mathfrak{U}},W,{\mathfrak{D}}^{\psi}_{al}\otimes A\right)\right]}.$ (24)
Proposition 6.3.
If a cycle ${\mathsf{f}}$ lies in the commutator DGLA (24), then $\psi(\langle{\mathsf{f}},\beta\rangle)$ depends only on the cohomology class of $\beta$ and the homology class of ${\mathsf{f}}$ in the DGLA (24).
Proof.
Suppose ${\mathsf{f}}\in C^{aug}_{p}({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A)$ and ${\mathsf{f}}=\partial{\mathsf{g}}$ for some ${\mathsf{g}}$ in the commutator DGLA. Then the vanishing of $\psi(\langle{\mathsf{f}},\beta\rangle)\in A$ follows from the proof of Lemma 6.1 and Lemma 6.2. Thus it remains to show that $\psi(\langle{\mathsf{f}},\check{\delta}b\rangle)=0$ for any $b\in C^{p-1}({\mathfrak{U}},\hat{W})$ if ${\mathsf{f}}$ is a cycle of the commutator DGLA. Indeed, since $\langle{\mathsf{f}},\partial^{*}b\rangle=\langle\partial{\mathsf{f}},b\rangle=0$, we have
$\langle{\mathsf{f}},\check{\delta}b\rangle=\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s}=-\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s},$
and the $\psi$-average of each term in the latter sum vanishes by Lemma 6.2. ∎
The evaluations of cycles of the commutator DGLA on spherical CS cocycles can be used to construct topological invariants of $G$-invariant gapped states on ${\mathbb{R}}^{n}$. We set $W={\mathbb{R}}^{n}$. The cycle we use as an input is the commutator class of $C_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ defined using the inhomogeneous Maurer-Cartan equation (Section 6.1). Since ${\mathfrak{D}}^{\psi,\bf G}_{al}$ is a sub-algebra of ${\mathfrak{D}}^{\psi}_{al}\otimes A$ with $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$, Prop. 6.3 applies to such cycles.
Let ${\mathfrak{U}}$ be a CS cover of ${\mathbb{R}}^{n}$, and $\beta$ be a spherical CS cocycle of $\hat{W}=S^{n-1}$ of degree $l$ with respect to the cover $\hat{\mathfrak{U}}$. Let ${\mathsf{p}}$ be a ${\mathsf{Q}}$-twisted MC element of $C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$. Evaluating $[{\mathsf{p}},{\mathsf{p}}]$ on $\beta$ and taking into account that the grading is shifted by $1$ relative to the Čech grading, we get a $G$-invariant element of ${{\mathfrak{d}}^{\psi}_{al}}\otimes{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$ where $2k-3=l$. Since $\psi$ is $G$-invariant, $\psi(\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ is a $G$-invariant element of ${\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$. Equivalently, it is the value of a degree $3$ linear function on cocycles of $\check{C}^{\bullet}(\hat{\mathfrak{U}},{\mathbb{R}})$ valued in ${\rm Sym}^{\bullet}{\mathfrak{g}}^{*}[2]$.
Theorem 6.1.
The function $\beta\mapsto\psi(\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ depends only on $\psi$ and the class of $({\mathfrak{U}},\beta)$ in the CS cohomology of $S^{n-1}$.
Proof.
For a fixed covering ${\mathfrak{U}}$, Prop. 6.1 and 6.3 implies $\psi(\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ is independent of ${\mathsf{p}}$. It depends on $\beta$ solely through its cohomology class. It remains to show invariance under refinement of cover.
Let $({\mathfrak{V}},\phi)$ be a refinement of ${\mathfrak{U}}$ with $V_{j}\subset U_{\phi(j)}.$ Cocycles $({\mathfrak{U}},\beta)$ and $({\mathfrak{V}},\phi^{*}\beta)$, where $(\phi^{*}\beta)_{j_{0},\dots,j_{k}}=\beta_{\phi(j_{0}),\dots,\phi(j_{k})}$, are in the same CS cohomology class. From Prop.6.1 there exists ${\mathsf{Q}}$-twisted MC element ${\mathsf{p}}$ for the cover ${\mathfrak{V}}$. Prop.6.2 then implies that $\phi_{*}{\mathsf{p}}$ is a ${\mathsf{Q}}$-twisted MC element ${\mathsf{p}}$ for the cover ${\mathfrak{U}}$. An easy expansion shows
$\langle[\phi_{*}{\mathsf{p}},\phi_{*}{\mathsf{p}}],\beta\rangle=\langle[{\mathsf{p}},{\mathsf{p}}],\phi^{*}\beta\rangle.$
Since any ${\mathsf{Q}}$-twisted MC element gives the same answer, we have shown that this contraction of interest depends only on the CS cohomology class of $({\mathfrak{U}},\beta)$. ∎
By Proposition 4.7, the cohomology group $H^{l}_{CS}(S^{n-1},{\mathbb{R}})$ is one-dimensioal for $l=n-1$ and vanishes otherwise. It is natural to take the class of $({\mathfrak{U}},\beta)$ to be a generator of $H^{n-1}_{CS}(S^{n-1},{\mathbb{Z}})$. This generator is uniquely defined once an orientation of ${\mathbb{R}}^{n-1}$ has been chosen. Thus for any even $n$ we obtain an invariant of gapped $G$-invariant states on ${\mathbb{R}}^{n}$ taking values in $G$-invariant polynomials on ${\mathfrak{g}}$ of degree $(n+2)/2$, and this invariant changes sign when the orientation of ${\mathbb{R}}^{n}$ is changed. This is in agreement with Chern-Simons field theory.
6.3 The Hall conductance
For physics applications, the most important case is $G=U(1)$ and $n=2$. Let us specialize the construction of topological invariants to this case.
Let $\psi$ be a $U(1)$-invariant gapped state of a lattice system on ${\mathbb{R}}^{n}$. The generator of the $U(1)$ action is ${\mathsf{Q}}\in{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Let ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$ be the sub-algebra of $U(1)$-invariant elements of ${\mathfrak{D}}^{\psi}_{al}$. Since $U(1)$ is connected, this the same as the sub-algebra of elements of ${\mathfrak{D}}_{al}$ which commute with ${\mathsf{Q}}$. More generally, for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}(U)={\mathfrak{D}}_{al}(U)\cap{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$. This is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. The graded local Lie algebra ${\mathfrak{D}}^{\psi,\bf G}_{al}$ reduces in this case to ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}\otimes{\mathbb{R}}[[t]]$ where $t$ is a variable of degree $-2$.
Pick a cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ of ${\mathbb{R}}^{n}$. To find a solution ${\mathsf{p}}$ of the inhomogenenous Maurer-Cartan equation with ${\mathsf{B}}={\mathsf{Q}}\otimes t$, we write ${\mathsf{p}}=\sum_{k=1}^{\infty}{\mathsf{p}}_{k}\otimes t^{k}$, where ${\mathsf{p}}_{k}\in C_{k-1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al})$. To compute the topological invariant of a state on ${\mathbb{R}}^{2}$ it is sufficient to solve for ${\mathsf{p}}_{1}$.
${\mathsf{p}}_{1}$ is a solution of $\partial{\mathsf{p}}_{1}={\mathsf{Q}}$. Explicitly, ${\mathsf{p}}_{1}=\{{\mathsf{Q}}_{i}\in{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}(U_{i})\}_{i\in I}$ such that $\sum_{i\in I}{\mathsf{Q}}_{i}={\mathsf{Q}}$. Such ${\mathsf{Q}}_{i}$ exist because ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$ is a cosheaf. Then the component of the commutator class in $C_{1}(U,{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al})$ is $\{[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\}_{i<j}$. The topological invariant of the state $\psi$ is obtained by evaluating it on a Čech 1-cocycle $\beta$ on $S^{1}$ corresponding to the cover $\hat{\mathfrak{U}}$ and then averaging the resulting inner derivation:
$\sigma(\beta)=\psi\left(\sum_{i<j}\beta_{ij}[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\right).$
Note that averaging over $\psi$ must be performed after the summation over $i,j$ because $[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]$ is not an inner derivation, in general.
The simplest CS cover of ${\mathbb{R}}^{2}$ which can represent a nontrivial class in $H^{1}_{CS}(S^{1},{\mathbb{R}})$ is made of three cones with a common vertex. In this case the construction of the invariant reduces to that in [7, 11]. It was shown there that the resulting invariant is proportional to the zero-temperature Hall conductance as determined by the Kubo formula.
6.4 Topological invariants of gapped states on subsets of ${\mathbb{R}}^{n}$
So far we assumed that $\Lambda$ is an arbitrary countable subset of ${\mathbb{R}}^{n}$ with a uniform $O(r^{n})$ bound on the number of points in a ball of radius $r$. Suppose $\Lambda\subset W^{\epsilon}$ for some CS set $W$ and some $\epsilon>0$. If this is the case, we will say the pair $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ describes a quantum lattice system on $W$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ the component ${\mathsf{F}}^{X}$ vanishes for any brick $X$ which does not intersect $W^{\epsilon}\cap\Lambda$, and thus for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ we have ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(U\wedge W)$ and ${\mathfrak{D}}^{\psi}_{al}(U)={\mathfrak{D}}^{\psi}_{al}(U\wedge W)$. Thus the assignment $U\mapsto{\mathfrak{D}}^{\psi}_{al}(U)$ can be regarded as a local Lie algebra over the site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$.
In particular, if a compact Lie group $G$ acts locally on the quantum lattice system on $W$ and preserves a gapped state $\psi$, the generator of the action ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ takes values in the sub-algebra ${\mathfrak{D}}^{\psi}_{al}(W)$. Consequently ${\mathfrak{D}}^{\psi,\bf G}_{al}$ is a graded local Fréchet-Lie algebra over ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ pointed by ${\mathsf{Q}}\in{\mathfrak{D}}^{\psi,\bf G}_{al}(W)$. Picking a cover ${\mathfrak{U}}$ of $W$ and applying the Čech functor gives an acyclic pointed DGLA whose commutator class can be evaluated on any CS cocycle $\beta$ of $\hat{W}$ to give an inner derivation. Its $\psi$-average is a $G$-invariant polynomial on ${\mathfrak{g}}$ which depends only on the class of $({\mathfrak{U}},\beta)$ in $H^{\bullet}(\hat{W},{\mathbb{R}})$. Thus a quantum lattice system on $W$ has a topological invariant which is a linear function of $H^{\bullet}(\hat{W},{\mathbb{R}})\rightarrow{\rm Sym}^{\bullet}{\mathfrak{g}}^{*}[2]$. It is easy to check that this linear function has degree $3$.
For example, consider a $U(1)$-invariant gapped state of a quantum lattice system on $W={\mathbb{R}}^{2}$ affinely embedded in ${\mathbb{R}}^{3}$. The topological invariant defined in Section 6.2 is obtained by contracting with a 2-cocycle of $S^{2}$ (the sphere at infinity) and vanishes for dimensional reasons. On the other hand, by contracting with the 1-cocycle of $\hat{W}=S^{1}$ one obtains the Hall conductance of this system.
For a more non-trivial example, for any $n>2$ consider a finite graph $\Gamma\subset S^{n-1}$ whose edges are geodesics and take $W$ to be the cone with base $\Gamma$ and apex at an arbitrary point of ${\mathbb{R}}^{n}$. The invariants of $U(1)$-invariant gapped states of quantum lattice systems on $W\subset{\mathbb{R}}^{n}$ are labeled by generators of the free abelian group $H^{1}(\Gamma,{\mathbb{Z}})$. This example goes beyond TQFT since $W$ need not be smooth or even locally Euclidean.
Appendix A 0-chains
In this section we use the results of [11] to derive the properties of LGAs which we used in Section 3.
A.1 0-chains on ${\mathbb{Z}}^{n}$
First let us characterize ${\mathfrak{D}}_{al}(U)$ in terms of 0-chains. A 0-chain on ${\mathbb{Z}}^{n}$ is an element ${\mathsf{a}}=\{{\mathsf{a}}_{j}\}_{j\in{\mathbb{Z}}^{n}}\in\prod_{j\in{\mathbb{Z}}^{n}}{\mathfrak{D}}_{al}(\{j\})$ such that
$\|{\mathsf{a}}\|_{k}:=\sup_{j\in{\mathbb{Z}}^{n}}\|{\mathsf{a}}_{j}\|_{\{j\},k}<\infty.$ (25)
We say a 0-chain ${\mathsf{a}}$ is supported on $U$ if ${\mathsf{a}}_{i}=0$ whenever $i\notin U$, and write $C_{0}(U)$ for the set of $U$-supported 0-chains, endowed with the norms (25) for $k\geq 0$.
Proposition A.1.
Let $U\subset{\mathbb{R}}^{n}$ be nonempty and let $U^{1}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq 1\}$ be its 1-thickening.
  1. i)
    If ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ then ${\mathsf{F}}=\partial{\mathsf{f}}$ for a $U^{1}$-supported 0-chain ${\mathsf{f}}$ with $\|{\mathsf{f}}\|_{k}\leq 2^{k}\|{\mathsf{F}}\|_{U,k}$
  2. ii)
    If ${\mathsf{f}}\in C_{0}(U)$, then for any $Y\in{\mathbb{B}}_{n}$ the sum
    $(\partial{\mathsf{f}})^{Y}:=\sum_{j\in{\mathbb{Z}}^{n}\cap U}{\mathsf{f}}_{j}^{Y}$
    is absolutely convergent and defines a map $\partial:C_{0}(U)\to{\mathfrak{D}}_{al}(U)$ with $\|\partial{\mathsf{f}}\|_{U,k}\leq C\|{\mathsf{f}}\|_{k+2n+1}$, where the constant $C>0$ depends only on $n$.
Lemma A.1.
For any nonempty $U,Y\subset{\mathbb{R}}^{n}$ we have
$1+d(Y,U^{1}\cap{\mathbb{Z}}^{n})+\operatorname{diam}(Y)\leq 2(1+d(Y,U)+\operatorname{diam}(Y))$
Proof.
Choose $y\in Y$ and $u\in U$ with $d(y,u)=d(Y,U)$, and choose $z\in{\mathbb{Z}}^{n}$ with $d(u,z)\leq 1$. Then since $z\in U^{1}\cap{\mathbb{Z}}^{n}$ we have
$d(Y,U^{1}\cap{\mathbb{Z}}^{n})$ $\leq\operatorname{diam}(Y)+d(y,z)$
$\leq\operatorname{diam}(Y)+d(y,u)+d(u,z)$
$\leq\operatorname{diam}(Y)+d(Y,U)+1,$
and the Lemma follows. ∎
Proof of Proposition A.1.
$i)$. Suppose ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$. Choose any total order on $U^{1}\cap{\mathbb{Z}}^{n}$ and for every $Y\in{\mathbb{B}}_{n}$ let $j^{*}(Y)$ be the closest point to $Y$ in $U^{1}\cap{\mathbb{Z}}^{n}$, using the total order as a tiebreaker. For every $i\in\Lambda$, define
${\mathsf{f}}_{i}:=\sum_{\begin{subarray}{c}Y\in{\mathbb{B}}_{n}\\ j^{*}(Y)=i\end{subarray}}{\mathsf{F}}^{Y}.$ (26)
Then either ${\mathsf{f}}_{i}^{Y}=0$ or $d(Y,U^{1}\cap{\mathbb{Z}}^{n})=d(Y,j)$ and $\|{\mathsf{f}}_{i}^{Y}\|=\|{\mathsf{F}}^{Y}\|$, and so
$\|{\mathsf{f}}_{i}\|_{k}$ $=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,\{i\})^{k}\|{\mathsf{f}}_{i}^{Y}\|$
$=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,U^{1}\cap{\mathbb{Z}}^{1}))^{k}\|{\mathsf{F}}^{Y}\|,$
which by Lemma A.1 is bounded by $2^{k}\|{\mathsf{F}}\|_{U,k}$.
$ii)$. Suppose that ${\mathsf{f}}$ is a $U$-supported 00-chain. For any $k\geq 0$ we have
$\|\partial{\mathsf{f}}^{Y}\|$ $\leq\sum_{j\in{\mathbb{Z}}^{n}\cap U}\|{\mathsf{f}}_{j}^{Y}\|$
$\leq\|{\mathsf{f}}\|_{k+2n+1}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-k-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$
It is not hard to show that for any brick $Y$ the quantity $(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$ is bounded by a constant $C$ depending only on $n$, which shows $\|\partial{\mathsf{f}}^{Y}\|\leq C\|{\mathsf{f}}\|_{k+2n+1}$. ∎
Proposition A.1 will allow us to apply the results of [11] on 0-chains. The results in [11] are phrased in terms of the norms
$\|{\mathsf{a}}\|^{KS}_{x,k}:=\sup_{r>0}(1+r)^{k}\inf_{{\mathsf{b}}\in{\mathfrak{d}}(B_{r}(x))}\|{\mathsf{a}}-{\mathsf{b}}\|$
where ${\mathfrak{d}}(B_{r}(x))$ is the set of traceless anti-hermitian operators strictly localized on the ball of radius $r$ around $x\in{\mathbb{R}}^{n}$. To apply their results we prove the equivalence of these norms:
Lemma A.2.
For any $x\in{\mathbb{Z}}^{n}$ and $k>0$, the norms $\|\cdot\|^{KS}_{x,k}$ and $\|\cdot\|_{\{x\},k}$ obey
$\|{\mathsf{a}}\|^{KS}_{x,k}$ $\leq C\|{\mathsf{a}}\|_{\{x\},k+2n+2}$ (27)
$\|{\mathsf{a}}\|_{\{x\},k}$ $\leq 8^{k}C^{\prime}\|{\mathsf{a}}\|^{KS}_{x,k}$ (28)
where $C,C^{\prime}$ are constants depending only on $k$.
Proof.
Suppose $\|{\mathsf{a}}\|_{\{x\},k+2n+1}<\infty$ and let $r>0$. Define ${\mathsf{b}}:=\sum_{X\subset B_{r}(x)}{\mathsf{a}}^{X}$. Then
$\|{\mathsf{b}}-{\mathsf{a}}\|\leq\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-k-2n-2}$
Since $X\nsubseteq B_{r}(x)$ means $\operatorname{diam}(X)+d(x,X)\geq r$, we continue this as follows
$\leq(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-2n-2}$
$\leq C(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}$
where in the last line we used Lemma 3.1. This proves (27). To prove (28) suppose $\|{\mathsf{a}}\|^{KS}_{x,k}<\infty$ and let $X$ be any brick. Set $r:=\lfloor(\operatorname{diam}(X)+d(x,X))/4\rfloor$. Then $X\nsubseteq B_{r}(x)$. Indeed, if $X\subseteq B_{r}(x)$ then $d(x,X)\leq r$ and $\operatorname{diam}(X)\leq 2r$, implying $\operatorname{diam}(X)+d(x,X)\leq 3r$, which is impossible. Choose ${\mathsf{b}}\in{\mathfrak{d}}(B_{r}(x))$ with $\|{\mathsf{a}}-{\mathsf{b}}\|\leq(1+r)^{-k}\|{\mathsf{a}}\|_{x,k}^{KS}$. Since $X\nsubseteq B_{r}(x)$ we have ${\mathsf{b}}^{X}=0$, and so
$\|{\mathsf{a}}^{X}\|$ $=\|({\mathsf{a}}-{\mathsf{b}})^{X}\|$
$\leq 4^{n}\|{\mathsf{a}}-{\mathsf{b}}\|$
$\leq 4^{n}\|{\mathsf{a}}\|_{x,k}^{KS}(1+r)^{-k}$
$\leq 4^{n+2k}\|{\mathsf{a}}\|_{x,k}^{KS}(1+\operatorname{diam}(X)+d(x,X))^{-k},$
where in the second line we used [11, Proposition C.1]. ∎
A.2 Proof of Lemma 3.7
[…]
Proof of Lemma 3.7.
$\mathcal{J}$ and $\mathcal{K}$ correspond to $\mathscr{J}_{{\mathsf{H}},w_{\Delta}}$ and $\mathscr{J}_{{\mathsf{H}},W_{\Delta}}$ in Section 4.1 of [11]. Part $i)$ follows from the definition of $\mathcal{K}$ and the fact that ${\mathsf{H}}$ preserves $\psi$. Part $ii)$ follows from Proposition A.1 and Lemma A.2, together with [11, Lemma F.1] (specifically line (177) therein). Part $iii)$ is [11, line (72)]. ∎
Appendix B Inhomogeneous Maurer-Cartan equation
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pointed pronilpotent DLGA which is a limit of an inverse system of nilpotent pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. The set ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ of ${\mathsf{B}}$-twisted MC elements has an additional equivalence relation. This section revolves around this additional structure leading eventually to a proof of Prop. 6.1.
We have morphisms $r_{N,K}:{\mathfrak{M}}_{N}\rightarrow{\mathfrak{M}}_{K}$ for all $N>K$ and the DGLA ${\mathfrak{M}}$ is the inverse limit of the corresponding system of DGLAs. For any $N\in{\mathbb{N}}$ let $r_{N}:{\mathfrak{M}}\rightarrow{\mathfrak{M}}_{N}$ be the natural projection. Let ${\mathfrak{j}}_{N}=\ker r_{N,N-1}$. The sets of ${\mathsf{B}}_{N}$-twisted MC elements of ${\mathfrak{M}}_{N}$ will be denoted ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$.
A ${\mathsf{B}}$-twisted MC element ${\mathsf{p}}$ gives rise to a degree $-1$ derivation $\partial_{\mathsf{p}}=\partial+{\rm ad}_{\mathsf{p}}$ of ${\mathfrak{M}}$ which squares to zero (twisted differential).
Lemma B.1.
If ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, then $r_{N,K}({\mathsf{p}})\in{\rm MC}({\mathfrak{M}}_{K},{\mathsf{B}}_{K})$ for all $K<N$. Further, ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ iff ${\mathsf{p}}^{N}=r_{N}({\mathsf{p}})\in{\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\ \forall N\in{\mathbb{N}}$.
Proof.
Straightforward. ∎
Thus ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is the inverse limit of the system of sets ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$.
We are going to define an equivalence relations on ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ for all $N$. This is done in the same way as for the ordinary (homogeneous) Maurer-Cartan equation [29, 30].
First, ${\mathfrak{M}}_{N,0}$ is a nilpotent Lie algebra, so there is a well-defined nilpotent Lie group $\exp({\mathfrak{M}}_{N,0})$ with the group law given by the Campbell-Baker-Hausdorff formula. Similarly, ${\mathfrak{M}}_{0}$ is pronilpotent (i.e. is an inverse limit of a system of nilpotent Lie algebras), so the CBH formula defines a group $\exp({\mathfrak{M}}_{0})$.
Second, there are Lie algebra homomorphisms from ${\mathfrak{M}}_{0}$ (resp. ${\mathfrak{M}}_{N,0}$) to the Lie algebras of affine-linear vector fields on ${\mathfrak{M}}_{-1}$ (resp. ${\mathfrak{M}}_{N,-1}$). This homomorphism maps ${\mathsf{a}}\in{\mathfrak{M}}_{0}$ or ${\mathfrak{M}}_{N,0}$ to the affine-linear vector field
$\xi_{\mathsf{a}}({\mathsf{p}})=[{\mathsf{a}},{\mathsf{p}}]-\partial{\mathsf{a}},$
where ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ or ${\mathfrak{M}}_{N,-1}$. Here we used the identification of the space of affine-linear vector fields on a vector space $V$ with the space of affine-linear maps $V\rightarrow V$. These homomorphisms exponentiate to actions of the groups $\exp({\mathfrak{M}}_{0})$ and $\exp({\mathfrak{M}}_{N,0})$ on ${\mathfrak{M}}_{-1}$ and ${\mathfrak{M}}_{N,-1}$ by affine-linear transformations. Explicitly, the actions are given by [29, 30]:
${\mathsf{p}}\mapsto\exp({\mathsf{a}})*{\mathsf{p}}=\exp({\rm ad}_{\mathsf{a}})({\mathsf{p}})+\frac{1-\exp({\rm ad}_{\mathsf{a}})}{{\rm ad}_{\mathsf{a}}}(\partial{\mathsf{a}}).$ (29)
Lemma B.2.
The actions of $\exp({\mathfrak{M}}_{0})$ on ${\mathfrak{M}}_{-1}$ (resp. $\exp({\mathfrak{M}}_{N,0})$ on ${\mathfrak{M}}_{N,-1}$) preserve the sets ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ (resp. ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$).
Proof.
The proof in [29], Section 1.3, applies just as well in the inhomogeneous case. ∎
We say that elements $p_{1},p_{2}$ of ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ or ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ are equivalent if they belong to the same orbit of these actions.
Remark B.1.
By analogy with the homogeneous case, one can define a ${\mathsf{B}}$-twisted Deligne groupoid as the transformation groupoid for the action of $\exp({\mathfrak{M}}_{0})$ on ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Similarly, one can define ”reduced” Deligne groupoids for every $N\in{\mathbb{N}}$.
We observe an easy but useful lemma.
Lemma B.3.
If ${\mathsf{p}}_{i}\in{\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $i=1,2$, are equivalent, then $r_{N,K}({\mathsf{p}}_{1})$ is equivalent to $r_{N,K}({\mathsf{p}}_{2})$ for all $K<N$.
Now come the interesting statements. Assume from now on that the DGLAs ${\mathfrak{M}}_{N}$ and ${\mathfrak{M}}$ are acyclic, that ${\mathfrak{j}}_{N}\subset{\mathfrak{M}}_{N}$ is central for all $N>1$, and that the morphisms $r_{N,N-1}$ are surjective for all $N>1$.
Lemma B.4.
With the above assumptions, the set ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty if and only if ${\rm MC}({\mathfrak{M}}_{1},{\mathsf{B}}_{1})$ is non-empty.
Proof.
The only if statement follows from Lemma B.1. To prove the if direction, we use induction on $N$. Assume ${\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$ is non-empty. Let ${\mathsf{p}}_{N-1}\in{\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$. Pick $\tilde{\mathsf{p}}_{N}\in r_{N,N-1}^{-1}({\mathsf{p}}_{N-1})$. Since ${\mathsf{p}}_{N-1}$ is a ${\mathsf{B}}_{N-1}$-twisted MC element and $r_{N,N-1}({\mathsf{B}}_{N})={\mathsf{B}}_{N-1}$, we must have
$\partial\tilde{\mathsf{p}}_{N}+\frac{1}{2}[\tilde{\mathsf{p}}_{N},\tilde{\mathsf{p}}_{N}]={\mathsf{B}}_{N}+{\mathsf{q}}_{N}$
for some ${\mathsf{q}}_{N}\in{\mathfrak{j}}_{N}$. We look for a solution of the ${\mathsf{B}}_{N}$-twisted MC equation of the form ${\mathsf{p}}_{N}=\tilde{\mathsf{p}}_{N}+{\mathsf{b}}_{N}$ where ${\mathsf{b}}_{N}\in{\mathfrak{j}}_{N}$. Taking into account that ${\mathfrak{j}}_{N}$ is central, the inhomogeneous MC equation reduces to $\partial{\mathsf{b}}_{N}=-{\mathsf{q}}_{N}$. Since by assumption ${\mathfrak{j}}_{N},{\mathfrak{M}}_{N}$, and ${\mathfrak{M}}_{N-1}$ form a short exact sequence and the latter two are acyclic, so is ${\mathfrak{j}}_{N}$. Hence such a ${\mathsf{b}}_{N}$ exists. This completes the inductive step proving that ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\neq\varnothing$ for all $N$. Moreover, we also proved that the morphisms ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\rightarrow{\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$ are surjective. Therefore by Lemma B.1 ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. ∎
The above lemma proves the first part of Prop. 6.1. Indeed when $[{\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$, ${\rm MC}({\mathfrak{M}}_{1},{\mathsf{B}}_{1})\neq\varnothing$ is implied by acyclicity. Next we show that with the above assumption on ${\mathfrak{M}}_{N}$ and ${\mathfrak{M}}$ all ${\mathsf{B}}$-twisted MC elements are equivalent.
Lemma B.5.
Suppose ${\mathsf{a}},{\mathsf{b}}\in{\mathfrak{M}}_{N,0}$ and ${\mathsf{a}}\in{\mathfrak{j}}_{N}$. Then for any ${\mathsf{p}}\in{\mathfrak{M}}_{N,-1}$ one has
$\exp({\mathsf{b}}+{\mathsf{a}})({\mathsf{p}})=\exp({\mathsf{b}})({\mathsf{p}})-\partial{\mathsf{a}}.$
Proof.
See [29], Lemma 2.8. ∎
Lemma B.6.
Let ${\mathsf{p}}_{i}\in{\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $i=1,2$ such that $r_{N,N-1}({\mathsf{p}}_{1})=r_{N,N-1}({\mathsf{p}}_{2})$. Then ${\mathsf{p}}_{1}$ and ${\mathsf{p}}_{2}$ are equivalent.
Proof.
Let ${\mathsf{q}}={\mathsf{p}}_{2}-{\mathsf{p}}_{1}\in{\mathfrak{M}}_{N,-1}$. By assumption, ${\mathsf{q}}\in{\mathfrak{j}}_{N}$. Moreover, $\partial{\mathsf{q}}=0$. Indeed:
$\partial{\mathsf{q}}=\frac{1}{2}[{\mathsf{p}}_{2},{\mathsf{p}}_{2}]-\frac{1}{2}[{\mathsf{p}}_{1},{\mathsf{p}}_{1}]=[{\mathsf{q}},{\mathsf{p}}_{1}]+\frac{1}{2}[{\mathsf{q}},{\mathsf{q}}]=0.$
By acyclicity of ${\mathfrak{j}}_{N}$, we have ${\mathsf{q}}=\partial{\mathsf{a}}$ for some ${\mathsf{a}}\in{\mathfrak{j}}_{N,0}$. Then Lemma B.5 implies that $\exp({\mathsf{a}})\in\exp({\mathfrak{M}}_{N,0})$ maps ${\mathsf{p}}_{2}$ to ${\mathsf{p}}_{1}$. ∎
Lemma B.7.
Let $\tilde{{\mathsf{p}}}_{i}\in{\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $i=1,2,$ be such that ${\mathsf{p}}_{1}=r_{N,N-1}(\tilde{{\mathsf{p}}}_{1})$ and ${\mathsf{p}}_{2}=r_{N,N-1}(\tilde{{\mathsf{p}}}_{2})$ are equivalent. Then $\tilde{{\mathsf{p}}}_{1}$ and $\tilde{{\mathsf{p}}}_{2}$ are equivalent.
Proof.
Let ${\mathsf{a}}\in{\mathfrak{M}}_{N-1,0}$ be an equivalence between ${\mathsf{p}}_{1}$ and ${\mathsf{p}}_{2}$, i.e. $\exp({\mathsf{a}})*{\mathsf{p}}_{2}={\mathsf{p}}_{1}.$ Let $\tilde{{\mathsf{a}}}\in{\mathfrak{M}}_{N,0}$ be any lift of ${\mathsf{a}}$. Then $r_{N,N-1}(\exp(\tilde{{\mathsf{a}}})*\tilde{{\mathsf{p}}}_{2})=r_{N,N-1}(\tilde{{\mathsf{p}}}_{1})$. By Lemma B.6, $\exp(\tilde{{\mathsf{a}}})*\tilde{{\mathsf{p}}}_{2}$ is equivalent to $\tilde{{\mathsf{p}}}_{1}$, therefore $\tilde{{\mathsf{p}}}_{2}$ is equivalent to $\tilde{{\mathsf{p}}}_{1}$. ∎
Proposition B.1.
For any $N\in{\mathbb{N}}$ all elements of ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ are equivalent. All elements of ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ are equivalent.
The first statement is proved by induction on $N$, where the inductive step is Lemma B.7. The second statement follows by passing to the inverse limit in $N$.
Theorem B.1.
Let ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ in $H_{\bullet}([{\mathfrak{M}},{\mathfrak{M}}])$ is independent of the choice of ${\mathsf{p}}$.
Proof.
Let ${\mathsf{p}},{\mathsf{p}}^{\prime}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and let ${\mathsf{a}}\in{\mathfrak{M}}_{0}$ be an equivalence between ${\mathsf{p}}$ and ${\mathsf{p}}^{\prime}$. Since ${\mathsf{p}},{\mathsf{p}}^{\prime}$ satisfy the inhomogeneous Maurer-Cartan equation and ${\mathsf{B}}$ is a cycle, $[{\mathsf{p}},{\mathsf{p}}]$ and $[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]$ are cycles as well. From (29) we have
${\mathsf{p}}^{\prime}-{\mathsf{p}}=-\partial{\mathsf{a}}+\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!},$
where
${\mathsf{f}}:=\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!}\in{\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}.$
Thus $[{\mathsf{p}},{\mathsf{p}}]-[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]=2\partial({\mathsf{p}}^{\prime}-{\mathsf{p}})=2\partial{\mathsf{f}}$ is $\partial$-exact in ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$. ∎